### v1

# Detecting and Rectifying Noisy Labels: A Similarity-based Approach

Abstract
Label noise in datasets could damage the performance of neural net training. As the size of modern deep networks grows, there is a growing demand for automated tools for detecting such errors. In this paper, we propose post-hoc, model-agnostic error detection and rectification methods utilizing the penultimate feature from a neural network. Our idea is based on the observation that the similarity between the penultimate feature of a mislabeled data point and its true class data points is higher than that for data points from other classes, making the probability of label occurrence within a tight, similar cluster informative for detecting and rectifying errors. Extensive experiments show our method not only demonstrates high performance across various noises but also automatically rectifies these errors to improve the quality of datasets [^1] Our implementation is available at https://github.com/huutiendang/noise-detection-and-rectification.
1 Introduction
While the majority of knowledge in AI systems is learned through unsupervised learning, supervised learning is an indispensable step in building strong AI systems (c.f. the LeCun’s cake). For Large Language Models (LLMs) such as GPTs Brown et al. (2020), LLaMA Touvron et al. (2023a, b), and Gemini Team et al. (2023), supervised learning accounts for only a small fraction of the total computation budget but has a significant impact on the models’ performance. Recent research (e.g. Zhou et al., 2023; Gunasekar et al., 2023) finds that high quality training data significantly improves performance while reducing the training cost by orders of magnitude. The need for automated tools for improving the quality of supervised learning data is rising as datasets and models are getting larger at an unprecedented speed.
Real-world datasets contain a notable amount of errors Beyer et al. (2020); Northcutt et al. (2021b). Previous works Dau et al. (2022); Nguyen-Duc et al. (2023) showed that removing errors from the training set improves the performance of AI models trained on that dataset. Automatic error rectification, however, is an underexplored topic. In this paper, we present a similarity based approach for error detection and rectification in large scale datasets. We provide a theoretical analysis from a similarity perspective, showing why the similarity between the penultimate feature of a mislabeled data point and its true class data points is often larger than that for data points from other classes (Sec. 3.2). Inspired by this observation, we develop simple yet effective similarity-based methods for detecting and rectifying label errors (Sec. 3.3). Extensive experiments demonstrate the superiority of our methods across various settings (Sec. 4.2). Furthermore, our methods are posthoc and model-agnostic i.e., they can be applied to any deep neural network (DNN) architectures without the need for retraining.
2 Background and Related Work
Notation.
[…]
Confident-based approach.
Confident-based methods are based on the notion of confident learning Northcutt et al. (2021a) that deriving label quality measurements by using predicted probability distribution Wang and Mueller (2022); Kuan and Mueller (2022); Thyagarajan et al. (2022). Low confidence serves as a heuristic indicating the likelihood of label noise. Given a data point $\mathbf{z}$ with output label $\mathbf{y}=(y_{1},...,y_{k},...y_{N})$, the model’s predicted probabilities is $\mathbf{p}=(p_{1},...,p_{k},...,p_{N})$ over $N$ classes. Northcutt et al. (2021a) proposed three label quality scoring methods:
(1) Self-Confidence (SC) refers to the estimated probability that the input $\mathbf{x}$ belongs to the class associated with its given label $k$: $\textnormal{SC}(\mathbf{z},\mathbf{p})=p_{y_{k}}$, for $k\in\{1,2,...,N\}$.
(2) Normalized-Margin (NM) is the quantified difference between the model’s estimated probability of the given label and the probability of the most likely class: $\textnormal{NM}(\mathbf{z},\mathbf{p})=p_{y_{k}}-p_{y_{j}{{}^{*}}}$, for $j^{*}=\operatorname*{arg\,max}_{j\neq k\in\{1,2,...,N\}}p_{y_{j}}$
(3) Confidence-Weighted Entropy (CE) is the ratio of SC score and the normalized entropy: $\textnormal{CE}(\mathbf{z},\mathbf{p})=\frac{p_{y_{k}}}{\mathcal{H}_{N}(\mathbf{p})}$, where $\mathcal{H}_{N}(\mathbf{p})=-\frac{1}{\textnormal{log}{N}}\displaystyle\sum_{n=1}^{N}p_{n}\log(p_{n})$.
Gradient-based approach.
Koh and Liang (2017) use Influence Function (IF)—a concept from robust statistics Hampel (1974)—for measuring the influence of a training data point on others in a DNN. Dau et al. (2022) proposed a way to adapt IF and its variants i.e., Gradient Dot Product (GD; Charpiat et al. (2019a)), Gradient Cosine (GC; Charpiat et al. (2019a)), and Tracing Gradient Decent (TracIn; Pruthi et al. (2020)), for identifying errors in large-scale source code datasets. The idea is that the gradients of error data points exhibit significantly large magnitudes and are opposite in direction to the gradients of normal data points. The algorithm computes the influence score of each data point in the noisy dataset with data points in a reference set. A more negative influence score means it is more likely to be an error. Nguyen-Duc et al. (2023) use class information to improve the performance and stability of these gradient methods.
(1) $\textnormal{IF}(\mathbf{z}^{(i)},\mathbf{z}^{(j)})=-\frac{1}{n}\mathbf{g}_{\hat{\theta}}(\mathbf{z}^{(i)})^{\top}\mathcal{H}_{\hat{\theta}}^{-1}\mathbf{g}_{\hat{\theta}}(\mathbf{z}^{(j)})$,
(2) $\textnormal{GD}(\mathbf{z}^{(i)},\mathbf{z}^{(j)})=\langle\mathbf{g}_{\hat{\theta}}(\mathbf{z}^{(i)}),\mathbf{g}_{\hat{\theta}}(\mathbf{z}^{(j)})\rangle$,
(3) $\textnormal{GC}(\mathbf{z}^{(i)},\mathbf{z}^{(j)})=\textnormal{cos}(\mathbf{g}_{\hat{\theta}}(\mathbf{z}^{(i)}),\mathbf{g}_{\hat{\theta}}(\mathbf{z}^{(j)}))$,
(4) $\textnormal{TracIn}(\mathbf{z}^{(i)},\mathbf{z}^{(j)})=\sum_{t=1}^{T}\eta_{t}\textnormal{GD}(\mathbf{z}^{(i)},\mathbf{z}^{(j)})$, where $\mathcal{H}_{\hat{\theta}}$ is the hessian matrix, $T$ is the number of epochs, and $\eta_{t}$ is the learning rate at epoch $t$.
Other approaches.
The rule-based approach Chu et al. (2013) and statistics-based approach Huang and He (2018) are commonly used for structured data such as tabular data. These methods are not suitable for deep learning, as they assume convexity in the model, and the rules in many large-scale datasets are not easy to find and describe.
3 Method
3.1 Observation
We design experiments to randomly corrupt and inject noise into datasets. We then train a deep network using gradient descent on these altered datasets to measure how noisy data points behave on other data points. As an illustrative example in Fig. 2(a), we observed that the similarity between the mislabeled data points and their true class data point penultimate-layer representations is often higher than that of other class data points.
(a) Distribution of (a) Cosine similarity and (b) Dot product over IMDB Maas et al. (2011) with $10\%$ noise. Blue bars represent the similarity between mislabeled data points and their true class data points, red bars represent the similarity between mislabeled data points and other class data points. Features are obtained from a trained BERT model.
We find that this phenomenon persists across varying percentages of noise. To complement our observation, we provide a theoretical analysis in Section 3.2.
3.2 Theoretical Analysis
We consider a deep network receives input $\mathbf{x}\in\mathbb{R}^{d}$, where $d$ is the input dimension. Let $\delta$ be the softmax activation function, $\mathbf{b}\in\mathbb{R}^{N}$ be the bias. Given an output function $\phi_{W}$ parameterized by $W\in\mathbb{R}^{N\times d}$. For two inputs $\mathbf{x}^{(i)}$ and $\mathbf{x}^{(j)}$ we have two output vectors $\phi_{W}(\mathbf{x}^{(i)})$ and $\phi_{W}(\mathbf{x}^{(j)})$ respectively. As seen by a deep network, we can measure the influence between $\mathbf{x}^{(i)}$ and $\mathbf{x}^{(j)}$ by quantifying how much $\phi_{W}(\mathbf{x}^{(i)})$ change would change $\phi_{W}(\mathbf{x}^{(j)})$ as well. If $\mathbf{x}^{(i)}$ and $\mathbf{x}^{(j)}$ have high similarity, then $\mathbf{x}^{(i)}$ have high influence on $\mathbf{x}^{(j)}$ and changing $\phi_{W}(\mathbf{x}^{(i)})$ have large effect on $\phi_{W}(\mathbf{x}^{(j)})$. Otherwise, if they have low similarity, then $\mathbf{x}^{(i)}$ have low influence on $\mathbf{x}^{(j)}$ and changing $\phi_{W}(\mathbf{x}^{(i)})$ have small effect on $\phi_{W}(\mathbf{x}^{(j)})$. To measure the similarity between data points, we employ a symmetric kernel proposed by Charpiat et al. (2019b): The Inner Product
$\mathcal{K}(\mathbf{x}^{(i)},\mathbf{x}^{(j)})=\left\langle\nabla_{W}\phi_{W}(\mathbf{x}^{(i)}),\nabla_{W}\phi_{W}(\mathbf{x}^{(j)})\right\rangle,$ (1)
We choice $\phi$ is the Cross Entropy between softmax activation output $\hat{\mathbf{y}}=\delta(W\mathbf{x}+\mathbf{b})$ and true distribution $\mathbf{y}$. We have $\phi_{W}(\mathbf{x})=\ell(\hat{\mathbf{y}},\mathbf{y};W)$. For simplicity, we remove the bias term. Let denote $\mathbf{u}=W\mathbf{x}$ and $\ell(\hat{\mathbf{y}},\mathbf{y})=\ell(\hat{\mathbf{y}},\mathbf{y};W)$. Using the chain rule:
$\nabla_{W}\ell(\hat{\mathbf{y}},\mathbf{y})$ $=\textnormal{vec}\left(\frac{\partial\ell(\hat{\mathbf{y}},\mathbf{y})}{\partial W}\right)$
$=\textnormal{vec}\left(\frac{\partial\ell(\hat{\mathbf{y}},\mathbf{y})}{\partial\mathbf{u}}\frac{\partial\mathbf{u}}{\partial W}\right)$
$=\nabla_{\mathbf{u}}\ell(\hat{\mathbf{y}},\mathbf{y})\mathbf{x}^{\top},$ (2)
[…]
$\frac{\partial\ell(\hat{\mathbf{y}},\mathbf{y})}{\partial\mathbf{u}}=\frac{\partial\ell(\hat{\mathbf{y}},\mathbf{y})}{\partial\hat{\mathbf{y}}}\frac{\partial\hat{\mathbf{y}}}{\partial\mathbf{u}}$ (3)
The first term on the right-hand side of Eqn. 3 is the partial derivatives of the loss w.r.t the predicted output $\hat{\mathbf{y}}$. Regarding to the fact: $\mathbf{y}$ is the one-hot vector present label $k$ has element $y_{k}=1$ and $y_{i}=0$ if $i\neq k$. We have:
$\frac{\partial\ell(\hat{\mathbf{y}},\mathbf{y})}{\partial\hat{\mathbf{y}}}$ $=\begin{bmatrix}\frac{\partial\ell(\hat{\mathbf{y}},\mathbf{y})}{\partial\hat{y}_{1}}\cdots\frac{\partial\ell(\hat{\mathbf{y}},\mathbf{y})}{\partial\hat{y}_{k}}\cdots\frac{\partial\ell(\hat{\mathbf{y}},\mathbf{y})}{\partial\hat{y}_{N}}\end{bmatrix}$
$=\begin{bmatrix}0\cdots\frac{1}{\hat{y}_{k}}\cdots 0\end{bmatrix}$ (4)
The second term is the matrix comprises partial derivatives of the predicted output $\hat{\mathbf{y}}$ w.r.t $\mathbf{u}$. Regarding to the fact:
$\frac{\partial\hat{\mathbf{y}}}{\partial\mathbf{u}}=\begin{bmatrix}\frac{\partial\hat{y}_{1}}{\partial u_{1}}&\frac{\partial\hat{y}_{1}}{\partial u_{2}}&\cdots&\frac{\partial\hat{y}_{1}}{\partial u_{N}}\\ \vdots&\vdots&\vdots&\vdots\\ \frac{\partial\hat{y}_{k}}{\partial u_{1}}&\frac{\partial\hat{y}_{k}}{\partial u_{2}}&\cdots&\frac{\partial\hat{y}_{k}}{\partial u_{N}}\\ \vdots&\vdots&\vdots&\vdots\\ \frac{\partial\hat{y}_{N}}{\partial u_{1}}&\frac{\partial\hat{y}_{N}}{\partial u_{2}}&\cdots&\frac{\partial\hat{y}_{N}}{\partial u_{N}}\\ \end{bmatrix}$ (5)
Substitute Eqn. 5 and Eqn. 4 into Eqn. 3, we get
$\frac{\partial\ell(\hat{\mathbf{y}},\mathbf{y})}{\partial\mathbf{u}}=$
$\begin{bmatrix}\frac{\partial\ell(\hat{\mathbf{y}},\mathbf{y})}{\partial\hat{y}_{k}}\frac{\partial\hat{y}_{k}}{\partial u_{1}}&\cdots&\frac{\partial\ell(\hat{\mathbf{y}},\mathbf{y})}{\partial\hat{y}_{k}}\frac{\partial\hat{y}_{k}}{\partial u_{k}}&\cdots&\frac{\partial\ell(\hat{\mathbf{y}},\mathbf{y})}{\partial\hat{y}_{k}}\frac{\partial\hat{y}_{k}}{\partial u_{N}}\end{bmatrix}$ (6)
Given a softmax activation function $\delta$ for class $k$: Case 1: $i=k$, we compute the derivative of softmax output $\hat{y}_{k}$ w.r.t $u_{k}$:
$\frac{\partial\hat{y}_{k}}{\partial u_{k}}$ $=\frac{\partial}{\partial u_{k}}\left(\frac{e^{u_{k}}}{\sum_{i=1}^{N}e^{u_{i}}}\right)$
$=\frac{e^{u_{k}}\left(\sum_{i=1}^{N}e^{u_{i}}\right)-e^{u_{k}}\cdot e^{u_{k}}}{\left(\sum_{i=1}^{N}e^{u_{i}}\right)^{2}}$
$=\hat{y}_{k}(1-\hat{y}_{k})$ (7)
Case 2: $i\neq k$, we compute the derivative of softmax output $\hat{y}_{k}$ w.r.t $u_{i}$:
$\frac{\partial\hat{y}_{k}}{\partial u_{i}}=\frac{\partial}{\partial u_{i}}\left(\frac{e^{u_{k}}}{\sum_{i=1}^{N}e^{u_{i}}}\right)$ (8)
Using the chain rule, we get:
$\frac{\partial\hat{y}_{k}}{\partial u_{i}}=-\frac{e^{u_{k}}\cdot e^{u_{i}}}{\left(\sum_{i=1}^{N}e^{u_{i}}\right)^{2}}=-\hat{y}_{k}\hat{y}_{i}$ (9)
Substitute Eqn. 7 and Eqn. 9 into Eqn. 6, we get a column vector:
$\nabla_{\mathbf{u}}\ell(\hat{\mathbf{y}},\mathbf{y})=\begin{bmatrix}-\hat{y}_{1}\cdots 1-\hat{y}_{k}\cdots-\hat{y}_{N}\end{bmatrix}^{\top}$ (10)
The inner product kernel in Eqn. 2 become:
$\mathcal{K}(\mathbf{x}^{(i)},\mathbf{x}^{(j)})=$ $\overbrace{\nabla_{\mathbf{u}}\ell(\hat{\mathbf{y}}^{(i)},\mathbf{y}^{(i)})^{\top}\nabla_{\mathbf{u}}\ell(\hat{\mathbf{y}}^{(j)},\mathbf{y}^{(j)})}^{\mathbf{G}^{(ij)}}$
$\cdot(\mathbf{x}^{(j)\top}\mathbf{x}^{(i)})$ (11)
We denote $\mathbf{G}^{(ij)}$ as the dot product of two gradients of the loss at $\mathbf{x}^{(i)}$ and $\mathbf{x}^{(j)}$. Suppose input $\mathbf{x}^{(i)}$ and input $\mathbf{x}^{(j)}$ have label $k$ and $k^{\prime}$ corresponding. We have
$\mathbf{G}^{(ij)}$ $=\begin{bmatrix}-\hat{y}^{(i)}_{1}\cdots 1-\hat{y}^{(i)}_{k}\cdots-\hat{y}^{(i)}_{N}\end{bmatrix}\begin{bmatrix}-\hat{y}^{(j)}_{1}\\ \vdots\\ 1-\hat{y}^{(j)}_{k^{\prime}}\\ \vdots\\ -\hat{y}^{(j)}_{N}\end{bmatrix}$ (12)
we consider 2 cases:
If $k=k^{\prime}$:
$\mathbf{G}^{(ij)}_{k=k^{\prime}}$ $=\hat{y}^{(i)}_{1}\hat{y}^{(j)}_{1}+\cdots+(1-\hat{y}^{(i)}_{k})(1-\hat{y}^{(j)}_{k})$
$+\cdots+\hat{y}^{(i)}_{N}\hat{y}^{(j)}_{N}$
$=(1-\hat{y}^{(i)}_{k})(1-\hat{y}^{(j)}_{k})+\sum_{n=1,n\neq k}^{N}\hat{y}^{(i)}_{n}\hat{y}^{(j)}_{n}$ (13)
If $k\neq k^{\prime}$:
$\mathbf{G}^{(ij)}_{k\neq{k^{\prime}}}$ $=\hat{y}^{(i)}_{1}\hat{y}^{(j)}_{1}+\cdots+(1-\hat{y}^{(i)}_{k})(-\hat{y}^{(j)}_{k})$+
$(1-\hat{y}^{(j)}_{k^{\prime}})(-\hat{y}^{(i)}_{k^{\prime}})+\cdots+\hat{y}^{(i)}_{N}\hat{y}^{(j)}_{N}$
$=\hat{y}^{(j)}_{k}(\hat{y}^{(i)}_{k}-1)+\hat{y}^{(i)}_{k^{\prime}}(\hat{y}^{(j)}_{k^{\prime}}-1)$
$+\sum_{n=1,n\neq k,n\neq k^{\prime}}^{N}\hat{y}^{(i)}_{n}\hat{y}^{(j)}_{n}$ (14)
During the training process, the model is more confident about the labels of data points, indicating the value of $\hat{y}^{(i)}_{k}$ and $\hat{y}^{(j)}_{k^{\prime}}$ being closer to $1$. Assume a well-trained model, and $\hat{y}^{(i)}_{k}\approx\hat{y}^{(j)}_{k^{\prime}}=\alpha$; $\hat{y}^{(i)}_{n}=\hat{y}^{(j)}_{n}\approx\epsilon=\frac{1-\alpha}{N-1}$. ($n\neq k$ and $n\neq k^{\prime}$). Substitute these values into Eqn. 13 and Eqn. 14, we get :
$\mathbf{G}^{(ij)}_{k=k^{\prime}}$ $\approx(1-\alpha)^{2}+\epsilon^{2}(N-1)$ (15)
$\mathbf{G}^{(ij)}_{k\neq{k^{\prime}}}$ $\approx-\frac{N(1-\alpha)^{2}}{(N-1)^{2}}=-\epsilon^{2}N$ (16)
[…]
$\frac{\mathcal{K}^{(ij)}_{k=k^{\prime}}}{\mathcal{K}^{(ij)}_{k\neq{k^{\prime}}}}=\frac{|\mathbf{G}^{(ij)}_{k=k^{\prime}}|}{|\mathbf{G}^{(ij)}_{k\neq{k^{\prime}}}|}$ $\approx\frac{|(1-\alpha)^{2}+\epsilon^{2}(N-1)|}{|-\epsilon^{2}N|}$
$\approx\frac{\epsilon^{2}(N-1)^{2}+\epsilon^{2}(N-1)}{\epsilon^{2}N}$
$\approx N-1$ (17)
[…]
3.3 Algorithm
[…]
Error Detection.
Given a noisy dataset $\mathcal{D}$, for each data point $\mathbf{z}^{(i)}\in\mathcal{D}$ with label $\mathbf{y}^{(i)}$, our algorithm finds $\mathcal{S}(\mathcal{D}_{\textnormal{aux}},\mathbf{z}^{(i)})$ such that every data point in $\mathcal{D}$ but not in $\mathcal{S}(\mathcal{D}_{\textnormal{aux}},\mathbf{z}^{(i)})$ is at most similar to $\mathbf{z}^{(i)}$ as the least similar point in $\mathcal{S}(\mathcal{D}_{\textnormal{aux}},\mathbf{z}^{(i)})$ (line 8). We define a scoring function that return $\mathbf{s}^{(i)}$—the probability of occurrence of label $\mathbf{y}^{(i)}$ in $\mathcal{S}(\mathcal{D}_{\textnormal{aux}},\mathbf{z}^{(i)})$ (line 9). The indicator $\mathbb{I}(\cdot)$ returns $1$ if the condition holds. The lower $\mathbf{s}$ is, the more likely a label error. We sort the data points in $\mathcal{D}$ in ascending order of $\mathbf{s}$ and obtain the sorted $\mathcal{D}^{\uparrow}$ (line 11).
Error Rectification.
We select the first $p\%$ samples of ranked set $\mathcal{D}^{\uparrow}$ denoted as $\mathcal{D}^{\uparrow}_{:p}$ and define a class decision rule $\textsc{Mode}(\cdot)$ that selects the label in $\mathcal{S}(\mathcal{D}_{\textnormal{aux}},\mathbf{z}^{(i)})$ has the highest probability and greater than threshold $\tau$. Otherwise, the label of $\mathbf{z}^{(i)}$ remains unchanged (line 13).
Algorithm 1 Error Detection and Rectification
1:
[…]
5:$k$: number of most similar data points.
6:noisy data points in $\mathcal{D}$ are rectified.
7:for $\mathbf{z}^{(i)}\in\mathcal{D}$ do
8:  $\mathcal{S}(\mathcal{D}_{\textnormal{aux}},\mathbf{z}^{(i)})=\{\mathbf{z}^{(j)}\in\mathcal{D}_{\textnormal{aux}}\}$ s.t. $|\mathcal{S}(\mathcal{D}_{\textnormal{aux}},\mathbf{z}^{(i)})|=k$, and
$\sigma(\mathbf{z}^{(i)},\mathbf{z}^{(j)})\geq\max_{\mathbf{z}^{\prime(i)}\in\mathcal{D}\backslash\mathcal{S}(\mathcal{D}_{\textnormal{aux}},\mathbf{z}^{(i)})}\sigma(\mathbf{z}^{(i)},\mathbf{z}^{\prime(i)})$
9:  $\mathbf{s}^{(i)}=\frac{1}{k}\sum_{\mathbf{z}^{(j)}\in\mathcal{S}(\mathcal{D}_{\textnormal{aux}},\mathbf{z}^{(i)})}\mathbb{I}(\mathbf{y}^{(j)}=\mathbf{y}^{(i)})$
10:end for
[…]
12:for $\mathbf{z}^{(i)}\in\mathcal{D}^{\uparrow}_{:p}$ do
13:  $\mathbf{z}^{(i)}=\left(\mathbf{x}^{(i)},\textsc{Mode}(\mathcal{S}(\mathcal{D}_{\textnormal{aux}},\mathbf{z}^{(i)}))\right)$
14:end for
15:return $\mathcal{D}$
4 Experiment
4.1 Experiment Setting
(a) Error detection accuracy of methods measure on Snippets. The x-axis in the figures presents the change of $t$ from $10\to 100\%$.
Table 1: Test accuracy after removing/rectifying potential noise samples on Snippets with 20% noise. The best and runner-up are marked.
Method Random noise Ambiguity noise Concentrated noise
Removed Rectified Removed Rectified Removed Rectified
Acc. (under noise) 88.64 (+0.00) 82.50 (+0.00) 79.38 (+0.00)
Confident-W. Entropy 83.02 (-5.62) 84.38 (+1.88) 77.50 (-1.88)
Normalize-Margin 87.19 (-1.45) 87.01 (+4.51) 77.89 (-1.49)
Self-Confidence 86.05 (-2.59) 87.36 (+4.86) 78.02 (-1.36)
Influence Function 83.24 (-5.40) 81.71 (-0.79) 79.51 (+0.13)
Gradient-Cosine 73.81 (-14.83) 83.33 (+0.83) 82.23 (+2.85)
Gradient-Dot 86.53 (-2.11) 82.01 (-0.49) 78.68 (-0.70)
TracIn 85.48 (-3.16) 81.71 (-0.79) 76.67 (-2.71)
Sim-Cos 89.43 (+0.79) 87.85 (-0.79) 85.00 (+2.50) 83.73 (+1.23) 81.53 (+2.15) 83.38 (+4.00)
Sim-Dot 86.71 (-1.93) 87.32 (-1.32) 82.98 (+0.48) 83.95 (+1.45) 81.53 (+2.15) 81.97 (+2.59)
Dataset and model.
We evaluate our method on two common benchmarks: Sentiment Analysis on IMDB Maas et al. (2011) and Topic Classification on Snippets Phan et al. (2008). We use BERT Devlin et al. (2019) as the standard model.
Modeling Realistic Noise.
We construct three realistic, human-originated types of noise: (1) Uniform noise: we randomly select data points and change the label to a different class. (2) Systematic ambiguity noise: we establish a rule $h$, which maps data points in a specific class to another fixed one. This means that the labels of selected instances in class $i$ are flipped to $h(i)$. To ensure distinctiveness, the mapping function $h$ adheres to the condition $h(i)\neq h(j)\forall i,j=\{1,...,N\}$, and $i\neq j$. This noise models the situations where inputs from multiple annotators are often aggregated; the resulting differences in annotations can serve as a model of systematic noise derived from human disagreements. (3) Concentrated noise: we select data points that are densely clustered and change their labels to target labels. We simulate scenarios where the datasets are poisoned by malicious to evaluate the sanitization ability of methods against data poisoning attacks.
Setting.
For each dataset, we construct groups of various sizes of noisy samples by corrupting the label of $p\%$ of the original training data. We construct the auxiliary dataset $\mathcal{D}_{\textnormal{aux}}$ by randomly selecting $m$ samples from the validation set. We fine-tune BERT on the noisy dataset $\mathcal{D}$ and select the best checkpoint measure on the validation set. We select the top $t\%$ ranked samples in $\mathcal{D}^{\uparrow}_{:p}$ and use error detection accuracy for evaluation. After rectifying/removing ranked samples (potentially noisy samples), we re-train the model and report the test accuracy and error reduction rate. Details of the dataset, model, and implementations are in Appendix A.
4.2 Main result
Error detection accuracy.
(1) Fig. 5(a) shows the error detection accuracy of methods with three types of noise with different percentages. As a result, when $t$ increases, the performance of gradient-based methods drastically decreases. This pattern is observed in all three types of noise across different percentages and in both Snippets (Fig. 9) and IMDB (Fig. 10). This result shows that the gradient-based methods are unstable and less inconsistent. (2) Confident-based methods are precise with uniform noises and systematic ambiguity noise, yet struggle with concentrated noise (Fig. 5(a)c). (3) Sim-Cos and Sim-Dot have high detection accuracy and slightly decrease when $t$ increases with different noise. This confirmed that the Similarity-based methods are effective and more robust to ambiguity and concentrated noises than gradient-/confident- based methods. (4) We observe that Sim-Dot often has low detection accuracy on IMDB (Fig. 10). We theoretically proved that for classification datasets with $N$ classes, the similarity of within-class data points is approximately N1N-1 times larger than that of other class data points. For IMDB, where $N=2$, this fraction becomes approximately $1$. That explains why Sim-Dot does not work well on IMDB. Mathematical details are in Section 3.2. (5) Sim-Cos consistently outperforms Sim-Dot by a large margin for all settings. We explain from the standpoint of feature normalization. By definition, Cosine similarity can be seen as the normalized Dot product. In Fig. 9 (b), we empirically show that the noisy samples have $L_{2}$-norm smaller than that of normal samples. Therefore, when dividing the feature of data points by their norm, the similarity between noisy and normal data points tends to be larger, leading to a more distinct distribution of similarities.
Improving datasets’ quality and model generalization.
Tab.1 shows a significant improvement in the test accuracy when removing/rectifying concentrated noise and systematic ambiguity noise. Nevertheless, a counterintuitive observation regarding random noise shows that removing/rectifying noise reduces the generalization of models, even when detection accuracy is high. We posit that deep models are robust to massive random noise Rolnick et al. (2017), then as the training process continues, the model also memorizes the noise, approaches the optimum, and the gradient of noise becomes smaller. The effect of noise, therefore, also decreases as the model converges. When removing noise, the model degrades the feature representation of noise samples and loses the generalization to unseen samples.
5 Conclusion
We introduce similarity-based algorithms for detecting and rectifying errors on large-scale datasets. We theoretically show that the similarity between the penultimate feature’s data points is useful for detecting errors. Experiment results demonstrated the superior performance of our methods, and their capability to improve the dataset quality and model generalization.
Limitations
We discuss the limitations of similarity-based methods: (1) The optimal detection accuracy of Sim-Cos and Sim-Dot unfortunately based on empirical validation, and depends on the choice of $k$ and $\mathcal{D}_{\textnormal{aux}}$ and also with different datasets and model architectures. (2) The generalization of models under the removal or rectification of noise remains uncertain, due to the limited exploration of datasets.
Ethics Statement
We consider only the public datasets and create artificial noises for evaluation. We do not pose any concern about the quality of the original datasets.
Appendix A Implementation detail
A.1 Dataset and model
Snippets Phan et al. (2008)
is an open dataset of web search snippets retrieved from Google Search with 8 domains including Business, Computers, Culture-Arts-Entertainment, Education-Science, Engineering, Health, Politics-Society, and Sports. The training data and testing data include $10,060$ and $2,280$ snippets respectively. For validation purposes, we randomly split original training data into train/valid with the ratio of $8048/2012$. The dataset can be found at http://jwebpro.sourceforge.net/data-web-snippets.tar.gz.
IMDB Maas et al. (2011)
is the most common benchmark for Sentiment Analysis task. IMDB includes $50,000$ reviews from the Internet Movie Database website with original $25,000$ negative and $25,000$ positive reviews. For validation purposes, we randomly split into training, validation, and test sets of sizes $15,000$, $5,000$, and $25,000$. The IMDB dataset can be found at https://ai.stanford.edu/˜amaas/data/sentiment/
BERT Devlin et al. (2019)
stands for Bidirectional Encoder Representations from Transformers. BERT is one of the most standard used pre-trained model for language understanding tasks. In all settings, we use BERT base uncased version.
A.2 Experiment detail
BERT was trained with AdamW Loshchilov and Hutter (2019) with learning rate $\eta=5e-5$, momentum $\beta=(0.9,0.999)$, cross entropy loss, batch-size of $16$ with $15$ epochs. For regularization, we use Dropout Srivastava et al. (2014) of $0.2$. We choose $p=\{5\%,10\%,20\%\}$, $m=1000$, $k=\{1,2,5,10,20,50,100,200\}$, $t=\{10\%,20\%,...,100\%\}$, and $\tau=0.8$. We compute the IF score for BERT with the last layer gradient as previous works Pezeshkpour et al. (2021); Hanawa et al. (2021) and use LiSSA Agarwal et al. (2017) to approximate the Hessian. For TracIn method, we calculate the influence score from the first epoch to the best epoch. We run experiments on $4$ seeds = $\{16,32,64,128\}$ and aggregate the results by taking the mean of these $4$ seeds. Two Nvidia A40s were used to run experiments.
Appendix B Additional results
B.1 Improving datasets and model generalization.
Tab.1 shows a significant improvement in the test accuracy when removing/rectifying concentrated noise and systematic ambiguity noise. Nevertheless, a counter-intuitive observation regarding random noise shows that removing/rectifying noise reduces the generalization of models, even when detection accuracy is high. We posit that deep models are robust to massive random noise Rolnick et al. (2017), then as the training process, the model also memorizes the noise, approaches the optimum, and the gradient of noise becomes smaller. The effect of noise, therefore, also decreases as the model converges. When removing noise, the model degrades the feature representation of noise samples and loses the generalization to unseen samples.
B.2 Effects of Hyperparameters
The effect of the size of $\mathcal{D}_{\textnormal{aux}}$.
We change the size of $\mathcal{D}_{\textnormal{aux}}$ from $100$ to $1500$, fix $k=100$. Tab. 2 and Tab. 3 show the change in error detection accuracy as the size of $\mathcal{D}_{\textnormal{aux}}$ changes. We observed that: (1) the detection accuracy of the methods increases as the size of $\mathcal{D}_{\textnormal{aux}}$ increases. This is true for both Snippets and IMDB, Sim-Cos and Sim-Dot, and for all levels of noise $\{5\%,10\%,20\%\}$. (2) The accuracy gains are larger for smaller $\mathcal{D}_{\textnormal{aux}}$ values. (3) Sim-Cos outperforms Sim-Dot for all noise levels and most $\mathcal{D}_{\textnormal{aux}}$ values.
Table 2: The effect of the size of $\mathcal{D}_{\textnormal{aux}}$, setting with Snippets.
Method 5% noise 10% noise 20% noise
Sim-Cos@100 $16.91$ $22.63$ $31.69$
Sim-Cos@200 $51.99$ $40.67$ $65.25$
Sim-Cos@300 $57.46$ $56.96$ $80.23$
Sim-Cos@400 $60.45$ $72.38$ $86.45$
Sim-Cos@500 $75.37$ $78.60$ $86.88$
Sim-Cos@600 $77.36$ $78.60$ $89.93$
Sim-Cos@700 $82.58$ $78.48$ $93.10$
Sim-Cos@800 $85.07$ $78.60$ $92.91$
Sim-Cos@900 $86.07$ $79.72$ $93.16$
Sim-Cos@1000 $86.56$ $80.01$ $93.28$
Sim-Cos@1500 $88.06$ $80.97$ $94.28$
Sim-Dot@100 $16.91$ $22.63$ $31.69$
Sim-Dot@200 $51.74$ $40.92$ $64.26$
Sim-Dot@300 $57.96$ $56.59$ $79.42$
Sim-Dot@400 $61.19$ $70.64$ $86.82$
Sim-Dot@500 $74.62$ $78.23$ $87.01$
Sim-Dot@600 $76.36$ $78.10$ $88.87$
Sim-Dot@700 $81.34$ $77.36$ $92.91$
Sim-Dot@800 $82.83$ $77.36$ $92.54$
Sim-Dot@900 $83.83$ $78.10$ $92.17$
Sim-Dot@1000 $84.08$ $78.73$ $92.41$
Sim-Dot@1500 $85.32$ $80.34$ $93.78$
Table 3: The effect of the size of $\mathcal{D}_{\textnormal{aux}}$ on detection performance with IMDB.
Method 5% noise 10% noise 20% noise
Sim-Cos@100 $6.40$ $8.55$ $20.10$
Sim-Cos@200 $58.10$ $74.00$ $78.55$
Sim-Cos@300 $60.30$ $75.60$ $78.55$
Sim-Cos@400 $60.10$ $75.65$ $78.55$
Sim-Cos@500 $60.20$ $75.55$ $78.55$
Sim-Cos@600 $59.90$ $75.70$ $78.52$
Sim-Cos@700 $60.20$ $75.55$ $78.45$
Sim-Cos@800 $60.10$ $75.60$ $78.47$
Sim-Cos@900 $60.20$ $75.60$ $78.42$
Sim-Cos@1000 $59.90$ $75.50$ $78.52$
Sim-Cos@1500 $60.10$ $75.55$ $78.50$
Sim-Dot@100 $6.40$ $8.55$ $20.10$
Sim-Dot@200 $57.80$ $66.65$ $77.72$
Sim-Dot@300 $58.00$ $65.70$ $77.80$
Sim-Dot@400 $58.10$ $68.20$ $78.47$
Sim-Dot@500 $57.90$ $67.10$ $77.95$
Sim-Dot@600 $58.00$ $73.65$ $78.35$
Sim-Dot@700 $58.50$ $69.45$ $78.52$
Sim-Dot@800 $57.90$ $65.70$ $77.87$
Sim-Dot@900 $58.20$ $68.65$ $77.47$
Sim-Dot@1000 $59.90$ $68.60$ $77.55$
Sim-Dot@1500 $57.90$ $68.05$ $78.10$
The effect of $k$.
Tab. 4 and Tab. 5 show that Sim-Cos and Sim-Dot have accuracy increase as $k$ increases for all noise levels $\{5\%,10\%,20\%\}$ and for most values of $k$. The accuracy gains are often larger for smaller values of $k$. Sim-Cos outperforms Sim-Dot, especially for larger values of $k$ for all noise levels and most values of $k$. The impact of noise on detection accuracy with Snippets is generally small, but it increases with $k$. For Sim-Cos with $20\%$ noise in Tab. 4, the accuracy starts to decrease after $k=20$. This suggests that using larger $k$ can harm the accuracy when the data is very noisy. For Sim-Dot with $10\%$ and $20\%$ noise in Tab. 4, the accuracy changes are less consistent across different values of $k$. This suggests that Sim-Dot may be more sensitive to the choice of $k$.
Table 4: Error detection accuracy of Sim-Cos and Sim-Dot changes as $k$ changes with Snippets.
Method 5% noise 10% noise 20% noise
Sim-Cos@k=1 $77.61$ $82.46$ $87.69$
Sim-Cos@k=2 $88.55$ $85.57$ $92.23$
Sim-Cos@k=5 $89.30$ $86.31$ $94.53$
Sim-Cos@k=10 $89.30$ $86.94$ $94.65$
Sim-Cos@k=20 $89.55$ $86.69$ $94.53$
Sim-Cos@k=50 $88.06$ $82.09$ $94.46$
Sim-Cos@k=100 $86.56$ $80.10$ $93.28$
Sim-Cos@k=200 $75.12$ $78.23$ $86.76$
Sim-Dot@k=1 $80.09$ $82.09$ $88.75$
Sim-Dot@k=2 $81.34$ $83.45$ $90.05$
Sim-Dot@k=5 $82.58$ $84.70$ $91.29$
Sim-Dot@k=10 $84.57$ $85.32$ $92.60$
Sim-Dot@k=20 $85.57$ $84.57$ $93.10$
Sim-Dot@k=50 $86.07$ $81.96$ $93.59$
Sim-Dot@k=100 $84.08$ $78.73$ $92.41$
Sim-Dot@k=200 $75.12$ $77.73$ $86.76$
Table 5: The effect of $k$ on detection performance with IMDB.
Method 5% noise 10% noise 20% noise
Sim-Cos@k=1 $28.50$ $50.85$ $66.15$
Sim-Cos@k=2 $49.10$ $66.80$ $69.72$
Sim-Cos@k=5 $57.80$ $72.05$ $75.65$
Sim-Cos@k=10 $59.50$ $74.10$ $76.77$
Sim-Cos@k=20 $59.50$ $74.90$ $77.42$
Sim-Cos@k=50 $60.20$ $75.55$ $78.42$
Sim-Cos@k=100 $59.90$ $75.50$ $78.52$
Sim-Cos@k=200 $60.00$ $75.45$ $78.57$
Sim-Dot@k=1 $57.50$ $67.25$ $75.17$
Sim-Dot@k=2 $57.00$ $67.30$ $75.30$
Sim-Dot@k=5 $57.20$ $67.95$ $75.80$
Sim-Dot@k=10 $57.10$ $67.75$ $76.00$
Sim-Dot@k=20 $57.10$ $68.15$ $76.27$
Sim-Dot@k=50 $57.90$ $68.55$ $76.95$
Sim-Dot@k=100 $59.90$ $68.60$ $77.55$
Sim-Dot@k=200 $58.20$ $68.25$ $78.40$
The effect of $\tau$ on error reduction rate.
We analyze the effect of $\tau$ on the error reduction rate. We vary the number of $\tau=\{0.5,0.6,0.7,0.8,0.9,0.99\}$. From results in Tab. 6 and Tab. 7, we see that for both Snippets and IMDB, the change of $\tau$ has minimal impact on the error reduction rate. Generally, the reduction rate is higher when the noise level is higher.
Table 6: Error Reduction Rate on Snippets.
Method 5% noise 10% noise 20% noise
[…]
Table 7: Error Reduction Rate on IMDB.
Method 5% noise 10% noise 20% noise
[…]
B.3 The correlation between detection methods.
We calculate the Spearman correlation between ranking scores assigned to samples by detection methods. Fig. 9(a)a shows that gradient-based, confident-based, and similarity-based methods have low Spearman correlation with each other. Confident-based and Gradient-based methods have negative Spearman correlation, indicating they are very different in ranking. We observed the same phenomenon in datasets across levels and types of noise.
(a) Visualization of (a) Spearman correlation, (b) the $L_{2}$-norm of noisy data points (presented by blue bars) and normal data points (presented by red bars). Using normalization (Cosine; (d)) enables the differentiation of the similarity distribution, therefore, enhancing detection accuracy compared to the unnormalized approach (Dot; (c)).
Table 8: Test accuracy on Snippets with 10% noise.
Method Random noise Ambiguity noise Concentrated noise
Removed Rectified Removed Rectified Removed Rectified
Acc. (under noise) 85.43 (+0.00) 88.24 (+0.00) 85.08 (+0.00)
Confidence-W. Entropy 88.07 (+2.64) 85.04 (-3.20) 84.56 (-0.52)
Normalize-Margin 81.75 (-3.68) 86.09 (-2.15) 79.73 (-5.35)
Self-Confidence 84.47 (-0.96) 87.06 (-1.18) 82.06 (-3.02)
Influence Function 90.04 (+4.61) 85.92 (-2.32) 81.36 (-3.72)
Gradient-Cosine 82.94 (-2.49) 81.18 (-7.06) 80.48 (-4.60)
Gradient-Dot 86.01 (+0.58) 85.13 (-3.11) 83.29 (-1.79)
TracIn 85.74 (+0.31) 86.88 (-1.36) 77.98 (-7.10)
Sim-Cos 79.73 (-5.70) 87.85 (+2.42) 85.48 (-2.76) 85.57 (-2.67) 81.58 (-3.23) 82.06 (-3.02)
Sim-Dot 88.29 (+2.86) 87.32 (+1.89) 84.34 (-3.90) 86.58 (-1.66) 81.62 (-3.46) 80.83 (-4.25)
Table 9: Test accuracy on Snippets with 5% noise.
Method Random noise Ambiguity noise Concentrated noise
Removed Rectified Removed Rectified Removed Rectified
Acc. (under noise) 87.14 (+0.00) 87.36 (+0.00) 87.93 (+0.00)
Confidence-W. Entropy 84.86 (-2.28) 86.57 (-0.79) 87.80 (-0.13)
Normalize-Margin 84.65 (-2.49) 84.82 (-2.54) 83.77 (-4.16)
Self-Confidence 85.35 (-1.79) 83.99 (-3.37) 86.09 (-1.84)
Influence Function 85.30 (-1.84) 86.00 (-1.36) 85.30 (-2.63)
Gradient-Cosine 85.43 (-1.71) 85.13 (-2.23) 79.51 (-8.42)
Gradient-Dot 84.56 (-2.58) 89.21 (+1.85) 87.06 (-0.87)
TracIn 84.65 (-2.49) 87.71 (+0.35) 81.88 (-6.05)
Sim-Cos 86.45 (-0.69) 85.17 (-1.97) 87.10 (-0.26) 85.48 (-1.88) 85.39 (-2.54) 87.19 (-0.74)
Sim-Dot 87.15 (+0.01) 87.06 (-0.08) 83.37 (-3.99) 84.34 (-3.02) 84.12 (-3.81) 79.82 (-8.11)
Table 10: Test accuracy on IMDB with 20% noise.
Method Random noise Ambiguity noise Concentrated noise
Removed Rectified Removed Rectified Removed Rectified
Acc. (under noise) 89.92 89.36 85.98
Confidence-W. Entropy 91.00 (+1.08) 90.73 (+1.37) 86.66 (+0.68)
Normalize-Margin 91.00 (+1.08) 90.73 (+1.37) 86.66 (+0.68)
Self-Confidence 91.00 (+1.08) 90.73 (+1.37) 86.66 (+0.68)
Influence Function 90.95 (+1.03) 90.51 (+1.15) 87.09 (+1.11)
Gradient-Cosine 90.59 (+0.67) 90.93 (+1.57) 86.78 (+0.80)
Gradient-Dot 90.43 (+0.51) 89.45 (+0.09) 82.87 (-3.20)
TracIn 90.19 (+0.27) 91.09 (+1.73) 82.87 (-3.20)
Sim-Cos 90.70 (+0.78) 88.73 (-1.19) 90.78 (+1.42) 90.83 (+1.47) 87.76 (+1.78) 84.73 (-1.25)
Sim-Dot 91.49 (+1.57) 90.46 (+0.54) 91.58 (+2.22) 90.13 (+0.77) 87.25 (+1.27) 87.46 (+1.48)
Table 11: Test accuracy on IMDB with 10% noise.
Method Random noise Ambiguity noise Concentrated noise
Removed Rectified Removed Rectified Removed Rectified
Acc. (under noise) 91.76 (+0.00) 91.68 (+0.00) 89.41 (+0.00)
Confidence-W. Entropy 91.79 (+0.03) 92.48 (+0.20) 90.95 (+1.54)
Normalize-Margin 91.79 (+0.03) 92.48 (+0.20) 90.95 (+1.54)
Self-Confidence 91.79 (+0.03) 92.48 (+0.20) 90.95 (+1.54)
Influence Function 91.65 (-0.11) 90.97 (-0.71) 86.26 (-3.15)
Gradient-Cosine 91.72 (-0.04) 91.29 (-0.39) 88.05 (-1.36)
Gradient-Dot 90.81 (-0.95) 92.10 (+0.42) 88.15 (-1.26)
TracIn 91.75 (-0.01) 91.66 (-0.02) 88.86 (-0.55)
Sim-Cos 91.12 (-0.64) 91.98 (+0.22) 92.52 (+0.84) 92.17 (+0.49) 90.15 (+0.74) 89.20 (-0.21)
Sim-Dot 91.38 (-0.38) 91.51 (-0.25) 91.52 (-0.16) 91.61 (-0.07) 89.78 (+0.37) 88.37 (-1.04)
Table 12: Test accuracy on IMDB with 5% noise.
Method Random noise Ambiguity noise Concentrated noise
Removed Rectified Removed Rectified Removed Rectified
Acc. (under noise) 91.67 (+0.00) 92.07 (+0.00) 91.52 (+0.00)
Confidence-W. Entropy 92.35 (+0.68) 92.58 (+0.51) 91.94 (+0.42)
Normalize-Margin 92.35 (+0.68) 92.58 (+0.51) 91.94 (+0.42)
Self-Confidence 92.35 (+0.68) 92.58 (+0.51) 91.94 (+0.42)
Influence Function 92.55 (+0.88) 91.76 (-0.31) 90.77 (-0.75)
Gradient-Cosine 92.46 (+0.79) 90.88 (-1.19) 89.78 (-1.74)
Gradient-Dot 91.79 (+0.12) 92.58 (+0.51) 90.32 (-1.20)
TracIn 92.62 (+0.95) 91.97 (-0.10) 91.45 (-0.07)
Sim-Cos 92.10 (+0.43) 92.10 (+0.43) 93.09 (+1.02) 92.50 (+0.43) 92.11 (+0.59) 91.74 (+0.22)
Sim-Dot 92.69 (+1.02) 92.06 (+0.39) 91.63 (-0.44) 91.09 (-0.98) 91.59 (+0.07) 90.96 (-0.56)
Figure 9: Detection accuracy of methods measured on Snippets with different levels of noise.
Figure 10: Detection accuracy of methods measured on IMDB with different levels of noise.

### v2

@@
 Abstract
-Label noise in datasets could damage the performance of neural net training.
-As the size of modern deep networks grows, there is a growing demand for automated tools for detecting such errors.
-In this paper, we propose post-hoc, model-agnostic error detection and rectification methods utilizing the penultimate feature from a neural network.
+Label noise in datasets could significantly damage the performance and robustness of deep neural networks (DNNs) trained on these datasets.
+As the size of modern DNNs grows, there is a growing demand for automated tools for detecting such errors.
+In this paper, we propose post-hoc, model-agnostic noise detection and rectification methods utilizing the penultimate feature from a DNN.
 Our idea is based on the observation that the similarity between the penultimate feature of a mislabeled data point and its true class data points is higher than that for data points from other classes, making the probability of label occurrence within a tight, similar cluster informative for detecting and rectifying errors.
-Extensive experiments show our method not only demonstrates high performance across various noises but also automatically rectifies these errors to improve the quality of datasets [^1] Our implementation is available at https://github.com/huutiendang/noise-detection-and-rectification.
-1 Introduction
-While the majority of knowledge in AI systems is learned through unsupervised learning, supervised learning is an indispensable step in building strong AI systems (c.f.
+Through theoretical and empirical analyses, we demonstrate that our approach achieves high detection performance across diverse, realistic noise scenarios and can automatically rectify these errors to improve dataset quality.
+Our implementation is available at https://anonymous.4open.science/r/noise-detection-and-rectification-AD8E.
+Keywords: Noisy label detection and rectification, post-hoc, model-agnostic, DNN, LLM
+[…]
+Detecting and Rectifying Noisy Labels:
+A Similarity-Based Approach
+Dang Huu-Tien, Minh-Phuong Nguyen, and Naoya Inoue†,‡
+JAIST,  RIKEN
+{tiendh, phuongnm, naoya-i}@jaist.ac.jp
+Abstract content
+1.  Introduction
+While the majority of knowledge in Artificial Intelligence (AI) systems is learned through unsupervised learning, supervised learning is an indispensable step in building strong AI systems (c.f.
 the LeCun’s cake).
-For Large Language Models (LLMs) such as GPTs Brown et al.
-(2020), LLaMA Touvron et al.
+For LLMs such as GPTs Brown et al.
+(2020b), Llama Touvron et al.
 (2023a, b), and Gemini Team et al.
 (2023), supervised learning accounts for only a small fraction of the total computation budget but has a significant impact on the models’ performance.
-Recent research (e.g.
-Zhou et al., 2023; Gunasekar et al., 2023) finds that high quality training data significantly improves performance while reducing the training cost by orders of magnitude.
-The need for automated tools for improving the quality of supervised learning data is rising as datasets and models are getting larger at an unprecedented speed.
-Real-world datasets contain a notable amount of errors Beyer et al.
+The efficacy and resilience of supervised models are susceptible to degradation due to noise in datasets.
+Noise (mislabeled) samples could result from mistakes in the labeling process or human disagreement Goh and Mueller (2023); Goh et al.
+(2022).
+Human error rates on real-world datasets could range from a few to tens of percent Beyer et al.
 (2020); Northcutt et al.
 (2021b).
+For more challenging datasets where expert knowledge is required, the error rate could be even higher.
+Recent research (e.g., Zhou et al., 2023; Gunasekar et al., 2023) finds that high-quality training data significantly improves performance while reducing the training cost by orders of magnitude.
+The large scale of modern datasets makes manual data inspection prohibitively expensive.
+The immense complexity of deep models makes it extremely hard to reliably and efficiently predict the effect a data point has on the performance of a deep model Basu et al.
+(2021).
+The need for automated tools for improving the quality of supervised learning data is rising as datasets and LLMs are getting larger at an unprecedented speed.
 Previous works Dau et al.
 (2022); Nguyen-Duc et al.
-(2023) showed that removing errors from the training set improves the performance of AI models trained on that dataset.
-Automatic error rectification, however, is an underexplored topic.
-In this paper, we present a similarity based approach for error detection and rectification in large scale datasets.
-We provide a theoretical analysis from a similarity perspective, showing why the similarity between the penultimate feature of a mislabeled data point and its true class data points is often larger than that for data points from other classes (Sec.
-3.2).
-Inspired by this observation, we develop simple yet effective similarity-based methods for detecting and rectifying label errors (Sec.
-3.3).
-Extensive experiments demonstrate the superiority of our methods across various settings (Sec.
-4.2).
-Furthermore, our methods are posthoc and model-agnostic i.e., they can be applied to any deep neural network (DNN) architectures without the need for retraining.
-2 Background and Related Work
+(2023) showed that removing noise from the training set improves the performance of AI models trained on that dataset.
+Automatic label noise rectification, however, is an underexplored topic.
+In this paper, we make the following contributions:
+  1. (1)
+    We first provide a theoretical analysis from a similarity perspective, showing why the similarity between the penultimate feature of a mislabeled data point and its true class data points is often larger than that for data points from other classes (Section 3.2).
+  2. (2)
+    Inspired by this observation, we develop a simple yet effective similarity-based approach with two representative methods for detecting and rectifying label noise in large-scale datasets (Section 3.3).
+Our methods are posthoc and model-agnostic, that is, they can be applied to any DNN architectures.
+  3. (3)
+    Extensive experiments demonstrate the superior performance of the proposed methods over gradient-based and confidence-based approaches, across diverse noise scenarios and scales (Section 5).
+2.  Background and Related Work
 Notation.
 […]
-Confident-based approach.
-Confident-based methods are based on the notion of confident learning Northcutt et al.
-(2021a) that deriving label quality measurements by using predicted probability distribution Wang and Mueller (2022); Kuan and Mueller (2022); Thyagarajan et al.
+2.1.  Confidence-Based Approach
+Confidence-based methods are based on the notion of confident learning Northcutt et al.
+(2021a) that derives label quality measurements by using predicted probability distribution Wang and Mueller (2022); Kuan and Mueller (2022); Thyagarajan et al.
 (2022).
@@
 (2021a) proposed three label quality scoring methods:
-(1) Self-Confidence (SC) refers to the estimated probability that the input $\mathbf{x}$ belongs to the class associated with its given label $k$: $\textnormal{SC}(\mathbf{z},\mathbf{p})=p_{y_{k}}$, for $k\in\{1,2,...,N\}$.
-(2) Normalized-Margin (NM) is the quantified difference between the model’s estimated probability of the given label and the probability of the most likely class: $\textnormal{NM}(\mathbf{z},\mathbf{p})=p_{y_{k}}-p_{y_{j}{{}^{*}}}$, for $j^{*}=\operatorname*{arg\,max}_{j\neq k\in\{1,2,...,N\}}p_{y_{j}}$
-(3) Confidence-Weighted Entropy (CE) is the ratio of SC score and the normalized entropy: $\textnormal{CE}(\mathbf{z},\mathbf{p})=\frac{p_{y_{k}}}{\mathcal{H}_{N}(\mathbf{p})}$, where $\mathcal{H}_{N}(\mathbf{p})=-\frac{1}{\textnormal{log}{N}}\displaystyle\sum_{n=1}^{N}p_{n}\log(p_{n})$.
-Gradient-based approach.
-Koh and Liang (2017) use Influence Function (IF)—a concept from robust statistics Hampel (1974)—for measuring the influence of a training data point on others in a DNN.
+Self-Confidence (SC) refers to the estimated probability that the input $\mathbf{x}$ belongs to the class associated with its given label $k$:
+$\textnormal{SC}(\mathbf{z},\mathbf{p})=p_{y_{k}};\;k\in\{1,2,...,N\}$ (1)
+Normalized-Margin (NM) is the quantified difference between the model’s estimated probability of the given label and the probability of the most likely class:
+$\textnormal{NM}(\mathbf{z},\mathbf{p})=p_{y_{k}}-p_{y_{j}{{}^{*}}};\;j^{*}=\operatorname*{arg\,max}_{j\neq k\in\{1,2,...,N\}}p_{y_{j}}$ (2)
+Confidence-Weighted Entropy (CE) is the ratio of SC score and the normalized entropy:
+$\textnormal{CE}(\mathbf{z},\mathbf{p})=\frac{p_{y_{k}}}{S_{N}(\mathbf{p})};S_{N}(\mathbf{p})=-\frac{1}{\textnormal{log}{N}}\displaystyle\sum_{n=1}^{N}p_{n}\log(p_{n})$ (3)
+$2.$  Gradient-Based Approach
+Koh and Liang (2017) use Influence Function (IF)—a concept from robust statistics Hampel (1974)—for measuring the influence of a training data point on others in a DNN without retraining.
+Several IF variants with faster measures have been proposed.
 Dau et al.
-(2022) proposed a way to adapt IF and its variants i.e., Gradient Dot Product (GD; Charpiat et al.
+(2022) proposed a way to adapt IF and its variants, that is, Gradient Dot Product (GD; Charpiat et al.
 (2019a)), Gradient Cosine (GC; Charpiat et al.
@@
 (2020)), for identifying errors in large-scale source code datasets.
-The idea is that the gradients of error data points exhibit significantly large magnitudes and are opposite in direction to the gradients of normal data points.
+The idea is that the gradients of noisy data points exhibit significantly large magnitudes and are opposite in direction to the gradients of normal data points.
 The algorithm computes the influence score of each data point in the noisy dataset with data points in a reference set.
-A more negative influence score means it is more likely to be an error.
+A more negative influence score means it is more likely to be a noisy data point.
 Nguyen-Duc et al.
 (2023) use class information to improve the performance and stability of these gradient methods.
-(1) $\textnormal{IF}(\mathbf{z}^{(i)},\mathbf{z}^{(j)})=-\frac{1}{n}\mathbf{g}_{\hat{\theta}}(\mathbf{z}^{(i)})^{\top}\mathcal{H}_{\hat{\theta}}^{-1}\mathbf{g}_{\hat{\theta}}(\mathbf{z}^{(j)})$,
-(2) $\textnormal{GD}(\mathbf{z}^{(i)},\mathbf{z}^{(j)})=\langle\mathbf{g}_{\hat{\theta}}(\mathbf{z}^{(i)}),\mathbf{g}_{\hat{\theta}}(\mathbf{z}^{(j)})\rangle$,
-(3) $\textnormal{GC}(\mathbf{z}^{(i)},\mathbf{z}^{(j)})=\textnormal{cos}(\mathbf{g}_{\hat{\theta}}(\mathbf{z}^{(i)}),\mathbf{g}_{\hat{\theta}}(\mathbf{z}^{(j)}))$,
-(4) $\textnormal{TracIn}(\mathbf{z}^{(i)},\mathbf{z}^{(j)})=\sum_{t=1}^{T}\eta_{t}\textnormal{GD}(\mathbf{z}^{(i)},\mathbf{z}^{(j)})$, where $\mathcal{H}_{\hat{\theta}}$ is the hessian matrix, $T$ is the number of epochs, and $\eta_{t}$ is the learning rate at epoch $t$.
-Other approaches.
+We summarize the gradient-based approach as follows:
+$\textnormal{IF}(\mathbf{z}^{(i)},\mathbf{z}^{(j)})$ $=-\frac{1}{n}\bm{g}_{\hat{\theta}}(\mathbf{z}^{(i)})^{\top}\bm{H}_{\hat{\bm{\theta}}}^{-1}\bm{g}_{\hat{\bm{\theta}}}(\mathbf{z}^{(j)})$ (4)
+$\textnormal{GD}(\mathbf{z}^{(i)},\mathbf{z}^{(j)})$ $=\left\langle\bm{g}_{\hat{\bm{\theta}}}(\mathbf{z}^{(i)}),\bm{g}_{\hat{\bm{\theta}}}(\mathbf{z}^{(j)})\right\rangle$ (5)
+[…]
+$\textnormal{TracIn}(\mathbf{z}^{(i)},\mathbf{z}^{(j)})$ $=\sum_{t=1}^{T}\eta_{t}\left\langle\bm{g}_{\bm{\theta}^{(t)}}(\mathbf{z}^{(i)}),\bm{g}_{{\bm{\theta}^{(t)}}}(\mathbf{z}^{(j)})\right\rangle$ (7)
+where $\bm{H}^{-1}_{\hat{\bm{\theta}}}$ is the inverse hessian matrix, $T$ is the number of epochs, and $\eta_{t}$ and $\bm{\theta}^{(t)}$ are the learning rate and optimal weights, at checkpoint $t$-th, respectively.
+$||\cdot||$ denote the Euclidean norm, $\langle\cdot,\cdot\rangle$ denote the inner product.
+2.3.  Other Approaches
 The rule-based approach Chu et al.
@@
 These methods are not suitable for deep learning, as they assume convexity in the model, and the rules in many large-scale datasets are not easy to find and describe.
-3 Method
-3.1 Observation
-We design experiments to randomly corrupt and inject noise into datasets.
-We then train a deep network using gradient descent on these altered datasets to measure how noisy data points behave on other data points.
-As an illustrative example in Fig.
-2(a), we observed that the similarity between the mislabeled data points and their true class data point penultimate-layer representations is often higher than that of other class data points.
+Bahri et al.
+(2020) use kkNN on the logit of a DNN to filter out noisy samples.
+Gao et al.
+(2016); Reeve and Kabán (2019) focus on examining the resilience of kkNN to random noise.
+Wu et al.
+(2020) proposed TopoFilter that filters noisy by utilizing the kkNN and Euclidean distance between pre-logits.
+Chong et al.
+(2022) uses the negative log probability of the assigned classes, akin to the Self-Confidence method.
+NoiseGPT Wang et al.
+(2024) leverages Multimodal LLMs to measure the “probability curvature” of data points.
+This curvature differs between benign and noisy samples.
+NoiseGPT then utilizes the In-Context Discrepancy to suggest or select corrected labels, serving as a knowledge expert for label noise detection and rectification.
+Unlike these prior works, this work proposes a post hoc, model-agnostic similarity-based approach for both detection and rectification, with theoretical justification under diverse, human-oriented noise.
+3.  Methodology
+3.1.  Motivation
+As a first start, we design a control experiment that randomly selects data points from the dataset and replaces their labels with random labels sampled from the output space.
 (a) Distribution of (a) Cosine similarity and (b) Dot product over IMDB Maas et al.
@@
 Blue bars represent the similarity between mislabeled data points and their true class data points, red bars represent the similarity between mislabeled data points and other class data points.
-Features are obtained from a trained BERT model.
-We find that this phenomenon persists across varying percentages of noise. To complement our observation, we provide a theoretical analysis in Section 3.2.
-3.2 Theoretical Analysis
-We consider a deep network receives input $\mathbf{x}\in\mathbb{R}^{d}$, where $d$ is the input dimension.
-Let $\delta$ be the softmax activation function, $\mathbf{b}\in\mathbb{R}^{N}$ be the bias.
-Given an output function $\phi_{W}$ parameterized by $W\in\mathbb{R}^{N\times d}$.
-For two inputs $\mathbf{x}^{(i)}$ and $\mathbf{x}^{(j)}$ we have two output vectors $\phi_{W}(\mathbf{x}^{(i)})$ and $\phi_{W}(\mathbf{x}^{(j)})$ respectively.
-As seen by a deep network, we can measure the influence between $\mathbf{x}^{(i)}$ and $\mathbf{x}^{(j)}$ by quantifying how much $\phi_{W}(\mathbf{x}^{(i)})$ change would change $\phi_{W}(\mathbf{x}^{(j)})$ as well.
-If $\mathbf{x}^{(i)}$ and $\mathbf{x}^{(j)}$ have high similarity, then $\mathbf{x}^{(i)}$ have high influence on $\mathbf{x}^{(j)}$ and changing $\phi_{W}(\mathbf{x}^{(i)})$ have large effect on $\phi_{W}(\mathbf{x}^{(j)})$.
-Otherwise, if they have low similarity, then $\mathbf{x}^{(i)}$ have low influence on $\mathbf{x}^{(j)}$ and changing $\phi_{W}(\mathbf{x}^{(i)})$ have small effect on $\phi_{W}(\mathbf{x}^{(j)})$.
+Features are obtained from a trained BERT.
+We then train a deep network using gradient descent on these altered datasets to measure how noisy data points behave on other data points in the penultimate latent space, i.e., the representation space before softmax.
+We then employ the cosine similarity and the dot product similarity to compute the similarity between the mislabeled data points and data points from other classes.
+Surprisingly, as an illustrative example in Fig.
+2(a), we observed that the similarity between the mislabeled data points and their true class data point penultimate-layer representations is often higher than that of other class data points.
+We hypothesize that in overparameterized models, when the noise rate is sufficiently small, around $5-10\%$—which is a realistic assumption in practice—the model tends to memorize the noise while approaching the optimum, causing the gradient contribution of the noisy samples to diminish.
+Consequently, as the model converges, the influence of noise on the decision boundary becomes negligible.
+To support this observation, we further provide a theoretical analysis in Section 3.2.
+3.2.  Theoretical Analysis
+Let us consider a DNN that receives input $\mathbf{x}\in\mathbb{R}^{d}$, where $d$ is the input dimension.
+Let $\delta$ be the softmax activation function, $\mathbf{b}\in\mathbb{R}^{N}$ be the bias term.
+Given a function $\phi_{\bm{W}}$ parameterized by $\bm{W}\in\mathbb{R}^{N\times d}$.
+For two inputs $\mathbf{x}^{(i)}$ and $\mathbf{x}^{(j)}$ we can obtain two output vectors $\phi_{\bm{W}}(\mathbf{x}^{(i)})$ and $\phi_{\bm{W}}(\mathbf{x}^{(j)})$ from $\phi_{\bm{W}}$, respectively.
+As seen by a DNN, we can measure the influence between $\mathbf{x}^{(i)}$ and $\mathbf{x}^{(j)}$ by quantifying how much $\phi_{\bm{W}}(\mathbf{x}^{(i)})$ change would change $\phi_{\bm{W}}(\mathbf{x}^{(j)})$ as well.
+If $\mathbf{x}^{(i)}$ and $\mathbf{x}^{(j)}$ have high similarity, then $\mathbf{x}^{(i)}$ have high influence on $\mathbf{x}^{(j)}$ and changing $\phi_{\bm{W}}(\mathbf{x}^{(i)})$ have large effect on $\phi_{\bm{W}}(\mathbf{x}^{(j)})$.
+Otherwise, if they have low similarity, then $\mathbf{x}^{(i)}$ have low influence on $\mathbf{x}^{(j)}$ and changing $\phi_{\bm{W}}(\mathbf{x}^{(i)})$ have small effect on $\phi_{\bm{W}}(\mathbf{x}^{(j)})$.
 To measure the similarity between data points, we employ a symmetric kernel proposed by Charpiat et al.
-(2019b): The Inner Product
-$\mathcal{K}(\mathbf{x}^{(i)},\mathbf{x}^{(j)})=\left\langle\nabla_{W}\phi_{W}(\mathbf{x}^{(i)}),\nabla_{W}\phi_{W}(\mathbf{x}^{(j)})\right\rangle,$ (1)
-We choice $\phi$ is the Cross Entropy between softmax activation output $\hat{\mathbf{y}}=\delta(W\mathbf{x}+\mathbf{b})$ and true distribution $\mathbf{y}$.
-We have $\phi_{W}(\mathbf{x})=\ell(\hat{\mathbf{y}},\mathbf{y};W)$.
+(2019b): The Inner Product kernel
+$\mathcal{K}(\mathbf{x}^{(i)},\mathbf{x}^{(j)})=\left\langle\nabla_{\bm{W}}\phi_{\bm{W}}(\mathbf{x}^{(i)}),\nabla_{\bm{W}}\phi_{\bm{W}}(\mathbf{x}^{(j)})\right\rangle,$ (8)
+We choice $\phi$ is the Cross Entropy between softmax activation output $\hat{\mathbf{y}}=\delta(\bm{W}\mathbf{x}+\mathbf{b})$ and true distribution $\mathbf{y}$.
+We have $\phi_{\bm{W}}(\mathbf{x})=\ell(\hat{\mathbf{y}},\mathbf{y};\bm{W})$.
 For simplicity, we remove the bias term.
-Let denote $\mathbf{u}=W\mathbf{x}$ and $\ell(\hat{\mathbf{y}},\mathbf{y})=\ell(\hat{\mathbf{y}},\mathbf{y};W)$.
-Using the chain rule:
-$\nabla_{W}\ell(\hat{\mathbf{y}},\mathbf{y})$ $=\textnormal{vec}\left(\frac{\partial\ell(\hat{\mathbf{y}},\mathbf{y})}{\partial W}\right)$
-$=\textnormal{vec}\left(\frac{\partial\ell(\hat{\mathbf{y}},\mathbf{y})}{\partial\mathbf{u}}\frac{\partial\mathbf{u}}{\partial W}\right)$
+Let denote $\mathbf{u}=\bm{W}\mathbf{x}$ and $\ell(\hat{\mathbf{y}},\mathbf{y})=\ell(\hat{\mathbf{y}},\mathbf{y};W)$.
+Using the chain rule, we obtain:
+$\nabla_{\bm{W}}\ell(\hat{\mathbf{y}},\mathbf{y})$ $=\textnormal{vec}\left[\frac{\partial\ell(\hat{\mathbf{y}},\mathbf{y})}{\partial\bm{W}}\right]=\textnormal{vec}\left[\frac{\partial\ell(\hat{\mathbf{y}},\mathbf{y})}{\partial\mathbf{u}}\frac{\partial\mathbf{u}}{\partial\bm{W}}\right]$
 $=\nabla_{\mathbf{u}}\ell(\hat{\mathbf{y}},\mathbf{y})\mathbf{x}^{\top},$ (9)
-[…]
+where vec[$\left[\frac{\partial\ell(\hat{\mathbf{y}},\mathbf{y})}{\partial\bm{W}}\right]$ is the vectorization of the derivative of the loss $\ell$ w.r.t.
+$\bm{W}$.
+The partial $\frac{\partial\ell(\hat{\mathbf{y}},\mathbf{y})}{\partial\mathbf{u}}$ is
 $\frac{\partial\ell(\hat{\mathbf{y}},\mathbf{y})}{\partial\mathbf{u}}=\frac{\partial\ell(\hat{\mathbf{y}},\mathbf{y})}{\partial\hat{\mathbf{y}}}\frac{\partial\hat{\mathbf{y}}}{\partial\mathbf{u}}$ (10)
 The first term on the right-hand side of Eqn.
-3 is the partial derivatives of the loss w.r.t the predicted output $\hat{\mathbf{y}}$.
+10 is the partial derivatives of the loss w.r.t.
+the predicted output $\hat{\mathbf{y}}$.
 Regarding to the fact: $\mathbf{y}$ is the one-hot vector present label $k$ has element $y_{k}=1$ and $y_{i}=0$ if $i\neq k$.
@@
 $=\begin{bmatrix}0\cdots\frac{1}{\hat{y}_{k}}\cdots 0\end{bmatrix}$ (11)
-The second term is the matrix comprises partial derivatives of the predicted output $\hat{\mathbf{y}}$ w.r.t $\mathbf{u}$. Regarding to the fact:
+The second term is the matrix that comprises partial derivatives of the predicted output $\hat{\mathbf{y}}$ w.r.t. $\mathbf{u}$. Regarding the fact:
 $\frac{\partial\hat{\mathbf{y}}}{\partial\mathbf{u}}=\begin{bmatrix}\frac{\partial\hat{y}_{1}}{\partial u_{1}}&\frac{\partial\hat{y}_{1}}{\partial u_{2}}&\cdots&\frac{\partial\hat{y}_{1}}{\partial u_{N}}\\ \vdots&\vdots&\vdots&\vdots\\ \frac{\partial\hat{y}_{k}}{\partial u_{1}}&\frac{\partial\hat{y}_{k}}{\partial u_{2}}&\cdots&\frac{\partial\hat{y}_{k}}{\partial u_{N}}\\ \vdots&\vdots&\vdots&\vdots\\ \frac{\partial\hat{y}_{N}}{\partial u_{1}}&\frac{\partial\hat{y}_{N}}{\partial u_{2}}&\cdots&\frac{\partial\hat{y}_{N}}{\partial u_{N}}\\ \end{bmatrix}$ (12)
-Substitute Eqn. 5 and Eqn. 4 into Eqn. 3, we get
+Substitute Eqn. 12 and Eqn. 11 into Eqn. 10, we get
 $\frac{\partial\ell(\hat{\mathbf{y}},\mathbf{y})}{\partial\mathbf{u}}=$
 $\begin{bmatrix}\frac{\partial\ell(\hat{\mathbf{y}},\mathbf{y})}{\partial\hat{y}_{k}}\frac{\partial\hat{y}_{k}}{\partial u_{1}}&\cdots&\frac{\partial\ell(\hat{\mathbf{y}},\mathbf{y})}{\partial\hat{y}_{k}}\frac{\partial\hat{y}_{k}}{\partial u_{k}}&\cdots&\frac{\partial\ell(\hat{\mathbf{y}},\mathbf{y})}{\partial\hat{y}_{k}}\frac{\partial\hat{y}_{k}}{\partial u_{N}}\end{bmatrix}$ (13)
-Given a softmax activation function $\delta$ for class $k$: Case 1: $i=k$, we compute the derivative of softmax output $\hat{y}_{k}$ w.r.t $u_{k}$:
+Given a softmax activation function $\delta$ for class $k$: Case 1: $i=k$, we compute the derivative of softmax output $\hat{y}_{k}$ w.r.t. $u_{k}$:
 $\frac{\partial\hat{y}_{k}}{\partial u_{k}}$ $=\frac{\partial}{\partial u_{k}}\left(\frac{e^{u_{k}}}{\sum_{i=1}^{N}e^{u_{i}}}\right)$
@@
 $=\hat{y}_{k}(1-\hat{y}_{k})$ (14)
-Case 2: $i\neq k$, we compute the derivative of softmax output $\hat{y}_{k}$ w.r.t $u_{i}$:
+Case 2: $i\neq k$, we compute the derivative of softmax output $\hat{y}_{k}$ w.r.t. $u_{i}$:
 $\frac{\partial\hat{y}_{k}}{\partial u_{i}}=\frac{\partial}{\partial u_{i}}\left(\frac{e^{u_{k}}}{\sum_{i=1}^{N}e^{u_{i}}}\right)$ (15)
-Using the chain rule, we get:
+Using the chain rule, we obtain:
 $\frac{\partial\hat{y}_{k}}{\partial u_{i}}=-\frac{e^{u_{k}}\cdot e^{u_{i}}}{\left(\sum_{i=1}^{N}e^{u_{i}}\right)^{2}}=-\hat{y}_{k}\hat{y}_{i}$ (16)
-Substitute Eqn. 7 and Eqn. 9 into Eqn. 6, we get a column vector:
+Substitute Eqn. 14 and Eqn. 16 into Eqn. 13, we get a column vector:
 $\nabla_{\mathbf{u}}\ell(\hat{\mathbf{y}},\mathbf{y})=\begin{bmatrix}-\hat{y}_{1}\cdots 1-\hat{y}_{k}\cdots-\hat{y}_{N}\end{bmatrix}^{\top}$ (17)
-The inner product kernel in Eqn. 2 become:
+The inner product kernel in Eqn. 8 becomes:
 $\mathcal{K}(\mathbf{x}^{(i)},\mathbf{x}^{(j)})=$ $\overbrace{\nabla_{\mathbf{u}}\ell(\hat{\mathbf{y}}^{(i)},\mathbf{y}^{(i)})^{\top}\nabla_{\mathbf{u}}\ell(\hat{\mathbf{y}}^{(j)},\mathbf{y}^{(j)})}^{\mathbf{G}^{(ij)}}$
 $\cdot(\mathbf{x}^{(j)\top}\mathbf{x}^{(i)})$ (18)
-We denote $\mathbf{G}^{(ij)}$ as the dot product of two gradients of the loss at $\mathbf{x}^{(i)}$ and $\mathbf{x}^{(j)}$.
+We denote $\mathbf{G}^{(ij)}$ as the product of two gradients of the loss at $\mathbf{x}^{(i)}$ and $\mathbf{x}^{(j)}$.
 Suppose input $\mathbf{x}^{(i)}$ and input $\mathbf{x}^{(j)}$ have label $k$ and $k^{\prime}$ corresponding.
@@
 20 and Eqn.
-14, we get :
+21, we get:
 $\mathbf{G}^{(ij)}_{k=k^{\prime}}$ $\approx(1-\alpha)^{2}+\epsilon^{2}(N-1)$ (22)
@@
 $\approx N-1$ (24)
+$\mathcal{K}^{(ij)}_{k=k^{\prime}}$ $\approx(N-1)\mathcal{K}^{(ij)}_{k\neq{k^{\prime}}}$ (25)
 […]
-3.3 Algorithm
+$3.$  Algorithm
 […]
-Error Detection.
-Given a noisy dataset $\mathcal{D}$, for each data point $\mathbf{z}^{(i)}\in\mathcal{D}$ with label $\mathbf{y}^{(i)}$, our algorithm finds $\mathcal{S}(\mathcal{D}_{\textnormal{aux}},\mathbf{z}^{(i)})$ such that every data point in $\mathcal{D}$ but not in $\mathcal{S}(\mathcal{D}_{\textnormal{aux}},\mathbf{z}^{(i)})$ is at most similar to $\mathbf{z}^{(i)}$ as the least similar point in $\mathcal{S}(\mathcal{D}_{\textnormal{aux}},\mathbf{z}^{(i)})$ (line 8).
-We define a scoring function that return $\mathbf{s}^{(i)}$—the probability of occurrence of label $\mathbf{y}^{(i)}$ in $\mathcal{S}(\mathcal{D}_{\textnormal{aux}},\mathbf{z}^{(i)})$ (line 9).
-The indicator $\mathbb{I}(\cdot)$ returns $1$ if the condition holds.
-The lower $\mathbf{s}$ is, the more likely a label error.
-We sort the data points in $\mathcal{D}$ in ascending order of $\mathbf{s}$ and obtain the sorted $\mathcal{D}^{\uparrow}$ (line 11).
-Error Rectification.
-We select the first $p\%$ samples of ranked set $\mathcal{D}^{\uparrow}$ denoted as $\mathcal{D}^{\uparrow}_{:p}$ and define a class decision rule $\textsc{Mode}(\cdot)$ that selects the label in $\mathcal{S}(\mathcal{D}_{\textnormal{aux}},\mathbf{z}^{(i)})$ has the highest probability and greater than threshold $\tau$.
-Otherwise, the label of $\mathbf{z}^{(i)}$ remains unchanged (line 13).
-Algorithm 1 Error Detection and Rectification
+Algorithm 1 Noise Detection and Rectification
 1:
@@
 15:return $\mathcal{D}$
-4 Experiment
-4.1 Experiment Setting
-(a) Error detection accuracy of methods measure on Snippets. The x-axis in the figures presents the change of $t$ from $10\to 100\%$.
-Table 1: Test accuracy after removing/rectifying potential noise samples on Snippets with 20% noise. The best and runner-up are marked.
+We define a scoring function that return $\mathbf{s}^{(i)}$—the probability of occurrence of label $\mathbf{y}^{(i)}$ in $\mathcal{S}(\mathcal{D}_{\textnormal{aux}},\mathbf{z}^{(i)})$ (line 9).
+The indicator $\mathbb{I}(\cdot)$ returns $1$ if the condition holds.
+The lower $\mathbf{s}$ is, the more likely a label noise.
+We sort the data points in $\mathcal{D}$ in ascending order of $\mathbf{s}$ and obtain the sorted $\mathcal{D}^{\uparrow}$ (line 11).
+We select the first $p\%$ samples of ranked set $\mathcal{D}^{\uparrow}$ denoted as $\mathcal{D}^{\uparrow}_{:p}$ and define a class decision rule $\textsc{Mode}(\cdot)$ that selects the label in $\mathcal{S}(\mathcal{D}_{\textnormal{aux}},\mathbf{z}^{(i)})$ has the highest probability and greater than threshold $\tau$.
+Otherwise, the label of $\mathbf{z}^{(i)}$ remains unchanged (line 13).
+4.  Experiment
+4.1.  Dataset and Model
+We use Snippets Phan et al.
+(2008) and IMDB Maas et al.
+(2011) for evaluation.
+Snippets is a dataset of web search snippets retrieved from Google Search with 8 domains, including business, computers, culture-arts-entertainment, education-science, engineering, health, politics-society, and sports.
+The training data and testing data include $10,060$ and $2,280$ snippets, respectively.
+For validation purposes, we randomly split the original training data into train/validation with the ratio of $8048/2012$.
+IMDB is the most common benchmark for sentiment analysis tasks.
+IMDB includes $50,000$ reviews from the Internet Movie Database website with original $25,000$ negative and $25,000$ positive reviews.
+For validation purposes, we randomly split into training, validation, and test sets of sizes $15,000$, $5,000$, and $25,000$.
+We use BERT Devlin et al.
+(2019) base uncased version, which is one of the most standard used models for language understanding tasks.
+We use open-weight LLMs: Llama-3.2-1B, Llama-3.2-3B, Llama-3.1-8B AI@Meta (2024) for the zero-shot detection scenarios in Subsection 5.2.
+Due to space constraints, we present the main results and include some additional results in the Appendix upon the final decision.
+4.2.  Experimental Setup
+(a) Noise detection accuracy of methods measured on Snippets (Fig. 8(a)a-c) and IMDB (Fig. 8(a)d-f) across various sizes and noise types.
+Table 1: Test accuracy of model measured on Snippets after removing or rectifying top $20\%$ potential noisy samples ranked by methods. The best, improvement, drop, and runner-up are marked.
 Method Random noise Ambiguity noise Concentrated noise
 Removed Rectified Removed Rectified Removed Rectified
-Acc. (under noise) 88.64 (+0.00) 82.50 (+0.00) 79.38 (+0.00)
-Confident-W. Entropy 83.02 (-5.62) 84.38 (+1.88) 77.50 (-1.88)
-Normalize-Margin 87.19 (-1.45) 87.01 (+4.51) 77.89 (-1.49)
-Self-Confidence 86.05 (-2.59) 87.36 (+4.86) 78.02 (-1.36)
-Influence Function 83.24 (-5.40) 81.71 (-0.79) 79.51 (+0.13)
-Gradient-Cosine 73.81 (-14.83) 83.33 (+0.83) 82.23 (+2.85)
-Gradient-Dot 86.53 (-2.11) 82.01 (-0.49) 78.68 (-0.70)
-TracIn 85.48 (-3.16) 81.71 (-0.79) 76.67 (-2.71)
-Sim-Cos 89.43 (+0.79) 87.85 (-0.79) 85.00 (+2.50) 83.73 (+1.23) 81.53 (+2.15) 83.38 (+4.00)
-Sim-Dot 86.71 (-1.93) 87.32 (-1.32) 82.98 (+0.48) 83.95 (+1.45) 81.53 (+2.15) 81.97 (+2.59)
-Dataset and model.
-We evaluate our method on two common benchmarks: Sentiment Analysis on IMDB Maas et al.
-(2011) and Topic Classification on Snippets Phan et al.
-(2008).
-We use BERT Devlin et al.
-(2019) as the standard model.
-Modeling Realistic Noise.
-We construct three realistic, human-originated types of noise: (1) Uniform noise: we randomly select data points and change the label to a different class.
-(2) Systematic ambiguity noise: we establish a rule $h$, which maps data points in a specific class to another fixed one.
-This means that the labels of selected instances in class $i$ are flipped to $h(i)$.
-To ensure distinctiveness, the mapping function $h$ adheres to the condition $h(i)\neq h(j)\forall i,j=\{1,...,N\}$, and $i\neq j$.
+Acc. (under noise) $88.64$ $82.50$ $79.38$
+Confidence-based Approach
+Confident-Weighted Entropy 83.0283.025.62-5.62 84.3884.38+1.88+1.88 77.5077.501.88-1.88
+Normalize-Margin $\underline{87.19}$1.45-1.45 $\underline{87.01}$+4.51+4.51 77.8977.891.49-1.49
+Self-Confidence 86.0586.052.59-2.59 87.36+4.86+4.86 78.0278.021.36-1.36
+Gradient-based Approach
+Influence Function 83.2483.245.40-5.40 81.7181.710.79-0.79 79.5179.51+0.13+0.13
+Gradient-Cosine 73.8173.8114.83-14.83 83.3383.33+0.83+0.83 82.23+2.85+2.85
+Gradient-Dot 86.5386.532.11-2.11 82.0182.010.49-0.49 78.6878.680.70-0.70
+TracIn 85.4885.483.16-3.16 81.7181.710.79-0.79 76.6776.672.71-2.71
+Similarity-based Approach
+Sim-Cos 89.43+0.79+0.79 87.8587.850.79-0.79 85.0085.00+2.50+2.50 $\underline{83.73}$+1.23+1.23 $\underline{81.53}$+2.15+2.15 83.3883.38+4.00+4.00
+Sim-Dot 86.7186.711.93-1.93 87.3287.321.32-1.32 82.9882.98+0.48+0.48 83.9583.95+1.45+1.45 81.5381.53+2.15+2.15 81.9781.97+2.59+2.59
+Table 2: Test accuracy of model measured on IMDB after removing or rectifying top $20\%$ potential noisy samples ranked by methods. The best, improvement, drops, and runner-up are marked.
+Method Random noise Ambiguity noise Concentrated noise
+Removed Rectified Removed Rectified Removed Rectified
+Acc. (under noise) $89.92$ $89.36$ $85.98$
+Confidence-based Approach
+Confidence-Weighted Entropy $\underline{91.00}$+1.08+1.08 90.7390.73+1.37+1.37 86.6686.66+0.68+0.68
+Normalize-Margin $\underline{91.00}$+1.08+1.08 90.7390.73+1.37+1.37 86.6686.66+0.68+0.68
+Self-Confidence $\underline{91.00}$+1.08+1.08 90.7390.73+1.37+1.37 86.6686.66+0.68+0.68
+Gradient-based Approach
+Influence Function 90.9590.95+1.03+1.03 90.5190.51+1.15+1.15 87.0987.09+1.11+1.11
+Gradient-Cosine 90.5990.59+0.67+0.67 90.9390.93+1.57+1.57 86.7886.78+0.80+0.80
+Gradient-Dot 90.4390.43+0.51+0.51 89.4589.45+0.09+0.09 82.8782.873.20-3.20
+TracIn 90.1990.19+0.27+0.27 $\underline{91.09}$+1.73+1.73 82.8782.873.20-3.20
+Similarity-based Approach
+Sim-Cos 90.7090.70+0.78+0.78 88.7388.731.19-1.19 90.7890.78+1.42+1.42 90.8390.83+1.47+1.47 87.76+1.78+1.78 84.7384.731.25-1.25
+Sim-Dot 91.49+1.57+1.57 90.4690.46+0.54+0.54 91.58+2.22+2.22 90.1390.13+0.77+0.77 $\underline{87.25}$+1.27+1.27 87.4687.46+1.48+1.48
+Modeling realistic noise. We construct three realistic, human-originated types of noise:
+(1) Random noise: we randomly select data points and change the label to a different class.
+(2) Systematic ambiguity noise: we establish a rule $r$, which maps data points in a specific class to another fixed one.
+This means that the labels of selected instances in class $i$ are flipped to $r(i)$.
+To ensure distinctiveness, the mapping function $r$ adheres to the condition $r(i)\neq r(j)\;\forall i,j=\{1,...,N\}$, and $i\neq j$.
 This noise models the situations where inputs from multiple annotators are often aggregated; the resulting differences in annotations can serve as a model of systematic noise derived from human disagreements.
 (3) Concentrated noise: we select data points that are densely clustered and change their labels to target labels.
-We simulate scenarios where the datasets are poisoned by malicious to evaluate the sanitization ability of methods against data poisoning attacks.
-Setting.
+We simulate scenarios where the datasets are poisoned by adversaries to evaluate the sanitization ability of methods against data poisoning attacks.
+Settings and hyperparameters.
 For each dataset, we construct groups of various sizes of noisy samples by corrupting the label of $p\%$ of the original training data.
@@
 We fine-tune BERT on the noisy dataset $\mathcal{D}$ and select the best checkpoint measure on the validation set.
-We select the top $t\%$ ranked samples in $\mathcal{D}^{\uparrow}_{:p}$ and use error detection accuracy for evaluation.
-After rectifying/removing ranked samples (potentially noisy samples), we re-train the model and report the test accuracy and error reduction rate.
-Details of the dataset, model, and implementations are in Appendix A.
-4.2 Main result
-Error detection accuracy.
-(1) Fig.
-5(a) shows the error detection accuracy of methods with three types of noise with different percentages.
-As a result, when $t$ increases, the performance of gradient-based methods drastically decreases.
-This pattern is observed in all three types of noise across different percentages and in both Snippets (Fig.
-9) and IMDB (Fig.
-10).
-This result shows that the gradient-based methods are unstable and less inconsistent.
-(2) Confident-based methods are precise with uniform noises and systematic ambiguity noise, yet struggle with concentrated noise (Fig.
-5(a)c).
-(3) Sim-Cos and Sim-Dot have high detection accuracy and slightly decrease when $t$ increases with different noise.
-This confirmed that the Similarity-based methods are effective and more robust to ambiguity and concentrated noises than gradient-/confident- based methods.
-(4) We observe that Sim-Dot often has low detection accuracy on IMDB (Fig.
-10).
-We theoretically proved that for classification datasets with $N$ classes, the similarity of within-class data points is approximately N1N-1 times larger than that of other class data points.
-For IMDB, where $N=2$, this fraction becomes approximately $1$.
-That explains why Sim-Dot does not work well on IMDB.
-Mathematical details are in Section 3.2.
-(5) Sim-Cos consistently outperforms Sim-Dot by a large margin for all settings.
-We explain from the standpoint of feature normalization.
-By definition, Cosine similarity can be seen as the normalized Dot product.
-In Fig.
-9 (b), we empirically show that the noisy samples have $L_{2}$-norm smaller than that of normal samples.
-Therefore, when dividing the feature of data points by their norm, the similarity between noisy and normal data points tends to be larger, leading to a more distinct distribution of similarities.
-Improving datasets’ quality and model generalization.
-Tab.1 shows a significant improvement in the test accuracy when removing/rectifying concentrated noise and systematic ambiguity noise.
-Nevertheless, a counterintuitive observation regarding random noise shows that removing/rectifying noise reduces the generalization of models, even when detection accuracy is high.
-We posit that deep models are robust to massive random noise Rolnick et al.
-(2017), then as the training process continues, the model also memorizes the noise, approaches the optimum, and the gradient of noise becomes smaller.
-The effect of noise, therefore, also decreases as the model converges.
-When removing noise, the model degrades the feature representation of noise samples and loses the generalization to unseen samples.
-5 Conclusion
-We introduce similarity-based algorithms for detecting and rectifying errors on large-scale datasets.
-We theoretically show that the similarity between the penultimate feature’s data points is useful for detecting errors.
-Experiment results demonstrated the superior performance of our methods, and their capability to improve the dataset quality and model generalization.
-Limitations
-We discuss the limitations of similarity-based methods: (1) The optimal detection accuracy of Sim-Cos and Sim-Dot unfortunately based on empirical validation, and depends on the choice of $k$ and $\mathcal{D}_{\textnormal{aux}}$ and also with different datasets and model architectures.
-(2) The generalization of models under the removal or rectification of noise remains uncertain, due to the limited exploration of datasets.
-Ethics Statement
-We consider only the public datasets and create artificial noises for evaluation. We do not pose any concern about the quality of the original datasets.
-Appendix A Implementation detail
-A.1 Dataset and model
-Snippets Phan et al. (2008)
-is an open dataset of web search snippets retrieved from Google Search with 8 domains including Business, Computers, Culture-Arts-Entertainment, Education-Science, Engineering, Health, Politics-Society, and Sports.
-The training data and testing data include $10,060$ and $2,280$ snippets respectively.
-For validation purposes, we randomly split original training data into train/valid with the ratio of $8048/2012$.
-The dataset can be found at http://jwebpro.sourceforge.net/data-web-snippets.tar.gz.
-IMDB Maas et al. (2011)
-is the most common benchmark for Sentiment Analysis task.
-IMDB includes $50,000$ reviews from the Internet Movie Database website with original $25,000$ negative and $25,000$ positive reviews.
-For validation purposes, we randomly split into training, validation, and test sets of sizes $15,000$, $5,000$, and $25,000$.
-The IMDB dataset can be found at https://ai.stanford.edu/˜amaas/data/sentiment/
-BERT Devlin et al. (2019)
-stands for Bidirectional Encoder Representations from Transformers.
-BERT is one of the most standard used pre-trained model for language understanding tasks.
-In all settings, we use BERT base uncased version.
-A.2 Experiment detail
-BERT was trained with AdamW Loshchilov and Hutter (2019) with learning rate $\eta=5e-5$, momentum $\beta=(0.9,0.999)$, cross entropy loss, batch-size of $16$ with $15$ epochs.
+We select the top $t\%$ ranked samples in $\mathcal{D}^{\uparrow}_{:p}$ and use detection accuracy for evaluation.
+After rectifying/removing ranked samples (potentially noisy samples), we re-train the model and report the test accuracy and noise reduction rate.
+Models were trained with AdamW Loshchilov and Hutter (2019) with learning rate 5e55e-5, momentum $\beta=(0.9,0.999)$, cross entropy loss, batch-size of $16$ with $15$ epochs.
 For regularization, we use Dropout Srivastava et al.
@@
 We choose $p=\{5\%,10\%,20\%\}$, $m=1000$, $k=\{1,2,5,10,20,50,100,200\}$, $t=\{10\%,20\%,...,100\%\}$, and $\tau=0.8$.
-We compute the IF score for BERT with the last layer gradient as previous works Pezeshkpour et al.
+We compute the IF score for BERT with the last layer gradient as suggested in previous works Pezeshkpour et al.
 (2021); Hanawa et al.
@@
 (2017) to approximate the Hessian.
-For TracIn method, we calculate the influence score from the first epoch to the best epoch.
-We run experiments on $4$ seeds = $\{16,32,64,128\}$ and aggregate the results by taking the mean of these $4$ seeds.
-Two Nvidia A40s were used to run experiments.
-Appendix B Additional results
-B.1 Improving datasets and model generalization.
-Tab.1 shows a significant improvement in the test accuracy when removing/rectifying concentrated noise and systematic ambiguity noise.
-Nevertheless, a counter-intuitive observation regarding random noise shows that removing/rectifying noise reduces the generalization of models, even when detection accuracy is high.
-We posit that deep models are robust to massive random noise Rolnick et al.
-(2017), then as the training process, the model also memorizes the noise, approaches the optimum, and the gradient of noise becomes smaller.
-The effect of noise, therefore, also decreases as the model converges.
-When removing noise, the model degrades the feature representation of noise samples and loses the generalization to unseen samples.
-B.2 Effects of Hyperparameters
-The effect of the size of $\mathcal{D}_{\textnormal{aux}}$.
-We change the size of $\mathcal{D}_{\textnormal{aux}}$ from $100$ to $1500$, fix $k=100$.
+For TracIn method, we calculate the influence score from the first checkpoint to the best checkpoint.
+We report results averaged over $4$ random seeds.
+Two Nvidia A40s, each 4545GB of GPU RAM, were used to run experiments.
+5.  Result and Analysis
+5.1.  Main Result
+Similarity-based methods demonstrate superior robustness and effectiveness.
+Fig.
+8(a) shows the detection accuracy of different methods under three types of noise at varying noise levels.
+As $t$ increases, the performance of gradient-based methods drops significantly, a pattern consistently observed across all noise types and percentages in both Snippets (Fig.
+8(a)a–c) and IMDB (Fig.
+8(a)d–f).
+These results indicate that gradient-based methods are unstable and inconsistent.
+In contrast, confidence-based methods perform well under random and systematic ambiguity noise but struggle with concentrated noise.
+Both Sim-Cos and Sim-Dot achieve high detection accuracy with only a slight degradation as $t$ increases, demonstrating that similarity-based methods are more effective and robust to ambiguity and concentrated noise than gradient-based and confidence-based approaches.
+Empirical validation for the claim in Section 3.2.
+We observe that Sim-Dot and Sim-Cos consistently have lower detection accuracy on IMDB than that of Snippets (Fig.
+8(a)).
+We theoretically proved that for classification datasets with $N$ classes, the similarity of within-class data points is approximately N1N-1 times larger than that of other class data points (c.f.
+Eqn.
+25).
+For IMDB, where $N=2$, this fraction becomes small.
+This explains why Sim-Dot and Sim-Cos achieve higher performance on Snippets than on IMDB.
+Why Sim-Cos consistently outperforms Sim-Dot?
+We explain from the standpoint of feature normalization.
+By definition, cosine similarity can be seen as the normalized dot product.
+We empirically observed that the noisy features’ norm is smaller than that of benign samples.
+Therefore, when dividing the feature of data points by their norm, the similarity between noisy and normal data points tends to be larger, leading to a more distinct distribution of similarities.
+Models’ performance on noise-removed or noise-rectified datasets.
+Tab.1 and Tab.
+2 show the test accuracy of BERT trained on Snippets and IMDB after removing or rectifying potential noisy samples.
+We observe that for more harmful types of noise—such as ambiguity or concentrated noise—removing or rectifying these samples improves model generalization.
+In contrast, a counterintuitive trend emerges for random noise: removing or rectifying such noise tends to reduce generalization, even when the noise detection accuracy is relatively high.
+This phenomenon further highlights the notion that random noise can act as a useful regularizer for enhancing models’ generalization, indicating that deep nets exhibit strong robustness even under random noise Rolnick et al.
+(2017).
+5.2.  Ablation Study
+Table 3: Detection accuracy of Sim-Cos and Sim-Dot across noise types at different levels using three LLMs on Snippets and IMDB.
+Model Random noise Ambiguity noise Concentrated noise
+$5\%$ Noise $10\%$ Noise $20\%$ Noise $5\%$ Noise $10\%$ Noise $20\%$ Noise $5\%$ Noise $10\%$ Noise $20\%$ Noise
+Cos Dot Cos Dot Cos Dot Cos Dot Cos Dot Cos Dot Cos Dot Cos Dot Cos Dot
+Snippets
+BERT (trained) $85.3$ $83.1$ $89.2$ $88.2$ $93.1$ $92.9$ $86.7$ $84.5$ $90.5$ $90.1$ $92.9$ $91.0$ $47.6$ $45.3$ $64.0$ $58.0$ $79.5$ $75.0$
+Llama-3.2-1B $45.0$ $36.3$ $57.2$ $47.8$ $66.8$ $60.4$ $42.7$ $40.2$ $61.0$ $57.9$ $72.3$ $61.6$ $38.2$ $34.2$ $58.0$ $47.8$ $72.8$ $62.6$
+Llama-3.2-3B $47.2$ $41.2$ $56.5$ $52.1$ $66.8$ $63.9$ $40.7$ $40.0$ $60.0$ $58.9$ $71.9$ $67.2$ $37.0$ $35.2$ $56.2$ $54.5$ $72.6$ $67.5$
+Llama-3.1-8B $51.2$ $50.7$ $61.9$ $60.3$ $71.4$ $68.0$ $43.2$ $41.2$ $65.0$ $63.8$ $75.4$ $69.4$ $41.7$ $40.5$ $63.1$ $57.7$ $79.8$ $73.2$
+IMDB
+BERT (trained) $65.0$ $59.5$ $73.2$ $68.0$ $78.8$ $76.9$ $67.8$ $64.2$ $73.0$ $70.1$ $78.8$ $78.5$ $55.6$ $45.0$ $63.2$ $58.2$ $60.0$ $52.0$
+Llama-3.2-1B $11.7$ $8.3$ $21.1$ $16.5$ $35.0$ $31.2$ $10.8$ $8.0$ $20.5$ $17.3$ $33.8$ $31.1$ $12.2$ $9.4$ $22.3$ $17.5$ $35.7$ $29.8$
+Llama-3.2-3B $21.4$ $12.2$ $27.9$ $21.1$ $39.3$ $34.2$ $21.9$ $11.5$ $25.8$ $21.0$ $38.1$ $33.6$ $20.9$ $12.3$ $29.7$ $21.5$ $40.2$ $33.6$
+Llama-3.1-8B $24.2$ $13.4$ $29.5$ $18.1$ $41.9$ $30.1$ $23.5$ $12.2$ $29.0$ $18.1$ $41.2$ $29.9$ $22.9$ $11.7$ $31.7$ $18.6$ $42.6$ $31.1$
+Effects of the size of auxiliary set.
+Figure 8: Detection accuracy of models on Snippets (left) and IMDB (right) across sizes of $\mathcal{D}_{\textnormal{aux}}$.
+Fig.
+8 shows the change in the detection accuracy as the size of $\mathcal{D}_{\textnormal{aux}}$ increases from $100$ to $1500$, fixing $k=100$.
+We observed that: the detection accuracy of the methods increases as the size of $\mathcal{D}_{\textnormal{aux}}$ increases for both Snippets and IMDB, Sim-Cos and Sim-Dot, and for all levels of noise.
+Sim-Cos outperforms Sim-Dot for all noise levels.
+Effects of $k$ and $\tau$.
+Figure 9: Detection accuracy of similarity-based methods on Snippets (left) and IMDB (right) for $k\in\{5,10,20,50,100,200\}$.
+Fig.
+9 shows that Sim-Cos consistently outperforms Sim-Dot as $k$ increases for all noise levels.
+The impact of $k$ on detection accuracy with Snippets is generally small; with performance beginning to decline beyond $k=20$.
+This suggests that using a larger $k$ may degrade the detection accuracy.
+For IMDB, the accuracy increases as $k$ increases, but quickly reaches a saturation point at $k$ around $20$.
+We analyze the effect of $\tau$ on the error reduction rate.
+Fig.
+10 shows that for both Snippets and IMDB, the change of $\tau$ has minimal impact on the error reduction rate.
+Generally, the reduction rate is higher when the noise level is higher.
+Figure 10: Noise reduction rate across threshold $\tau\in\{0.5,0.6,0.7,0.8,0.9,0.99\}$.
+Generalization to LLMs.
+The size of the model and the diversity of its pretraining data endow LLM with strong capabilities on downstream tasks, even without explicit task-specific fine-tuning Brown et al.
+(2020a).
+One might ask: “Given a general-purpose LLM, which has not been trained on noisy datasets, how effectively can the proposed methods perform?” We aim to evaluate the effectiveness of similarity-based methods with three open-weight LLMs, including Llama-3.2-1B, Llama-3.2-3B, and Llama-3.1-8B.
 Tab.
-2 and Tab.
-3 show the change in error detection accuracy as the size of $\mathcal{D}_{\textnormal{aux}}$ changes.
-We observed that: (1) the detection accuracy of the methods increases as the size of $\mathcal{D}_{\textnormal{aux}}$ increases.
-This is true for both Snippets and IMDB, Sim-Cos and Sim-Dot, and for all levels of noise $\{5\%,10\%,20\%\}$.
-(2) The accuracy gains are larger for smaller $\mathcal{D}_{\textnormal{aux}}$ values.
-(3) Sim-Cos outperforms Sim-Dot for all noise levels and most $\mathcal{D}_{\textnormal{aux}}$ values.
-Table 2: The effect of the size of $\mathcal{D}_{\textnormal{aux}}$, setting with Snippets.
-Method 5% noise 10% noise 20% noise
-Sim-Cos@100 $16.91$ $22.63$ $31.69$
-Sim-Cos@200 $51.99$ $40.67$ $65.25$
-Sim-Cos@300 $57.46$ $56.96$ $80.23$
-Sim-Cos@400 $60.45$ $72.38$ $86.45$
-Sim-Cos@500 $75.37$ $78.60$ $86.88$
-Sim-Cos@600 $77.36$ $78.60$ $89.93$
-Sim-Cos@700 $82.58$ $78.48$ $93.10$
-Sim-Cos@800 $85.07$ $78.60$ $92.91$
-Sim-Cos@900 $86.07$ $79.72$ $93.16$
-Sim-Cos@1000 $86.56$ $80.01$ $93.28$
-Sim-Cos@1500 $88.06$ $80.97$ $94.28$
-Sim-Dot@100 $16.91$ $22.63$ $31.69$
-Sim-Dot@200 $51.74$ $40.92$ $64.26$
-Sim-Dot@300 $57.96$ $56.59$ $79.42$
-Sim-Dot@400 $61.19$ $70.64$ $86.82$
-Sim-Dot@500 $74.62$ $78.23$ $87.01$
-Sim-Dot@600 $76.36$ $78.10$ $88.87$
-Sim-Dot@700 $81.34$ $77.36$ $92.91$
-Sim-Dot@800 $82.83$ $77.36$ $92.54$
-Sim-Dot@900 $83.83$ $78.10$ $92.17$
-Sim-Dot@1000 $84.08$ $78.73$ $92.41$
-Sim-Dot@1500 $85.32$ $80.34$ $93.78$
-Table 3: The effect of the size of $\mathcal{D}_{\textnormal{aux}}$ on detection performance with IMDB.
-Method 5% noise 10% noise 20% noise
-Sim-Cos@100 $6.40$ $8.55$ $20.10$
-Sim-Cos@200 $58.10$ $74.00$ $78.55$
-Sim-Cos@300 $60.30$ $75.60$ $78.55$
-Sim-Cos@400 $60.10$ $75.65$ $78.55$
-Sim-Cos@500 $60.20$ $75.55$ $78.55$
-Sim-Cos@600 $59.90$ $75.70$ $78.52$
-Sim-Cos@700 $60.20$ $75.55$ $78.45$
-Sim-Cos@800 $60.10$ $75.60$ $78.47$
-Sim-Cos@900 $60.20$ $75.60$ $78.42$
-Sim-Cos@1000 $59.90$ $75.50$ $78.52$
-Sim-Cos@1500 $60.10$ $75.55$ $78.50$
-Sim-Dot@100 $6.40$ $8.55$ $20.10$
-Sim-Dot@200 $57.80$ $66.65$ $77.72$
-Sim-Dot@300 $58.00$ $65.70$ $77.80$
-Sim-Dot@400 $58.10$ $68.20$ $78.47$
-Sim-Dot@500 $57.90$ $67.10$ $77.95$
-Sim-Dot@600 $58.00$ $73.65$ $78.35$
-Sim-Dot@700 $58.50$ $69.45$ $78.52$
-Sim-Dot@800 $57.90$ $65.70$ $77.87$
-Sim-Dot@900 $58.20$ $68.65$ $77.47$
-Sim-Dot@1000 $59.90$ $68.60$ $77.55$
-Sim-Dot@1500 $57.90$ $68.05$ $78.10$
-The effect of $k$.
-Tab.
-4 and Tab.
-5 show that Sim-Cos and Sim-Dot have accuracy increase as $k$ increases for all noise levels $\{5\%,10\%,20\%\}$ and for most values of $k$.
-The accuracy gains are often larger for smaller values of $k$.
-Sim-Cos outperforms Sim-Dot, especially for larger values of $k$ for all noise levels and most values of $k$.
-The impact of noise on detection accuracy with Snippets is generally small, but it increases with $k$.
-For Sim-Cos with $20\%$ noise in Tab.
-4, the accuracy starts to decrease after $k=20$.
-This suggests that using larger $k$ can harm the accuracy when the data is very noisy.
-For Sim-Dot with $10\%$ and $20\%$ noise in Tab.
-4, the accuracy changes are less consistent across different values of $k$.
-This suggests that Sim-Dot may be more sensitive to the choice of $k$.
-Table 4: Error detection accuracy of Sim-Cos and Sim-Dot changes as $k$ changes with Snippets.
-Method 5% noise 10% noise 20% noise
-Sim-Cos@k=1 $77.61$ $82.46$ $87.69$
-Sim-Cos@k=2 $88.55$ $85.57$ $92.23$
-Sim-Cos@k=5 $89.30$ $86.31$ $94.53$
-Sim-Cos@k=10 $89.30$ $86.94$ $94.65$
-Sim-Cos@k=20 $89.55$ $86.69$ $94.53$
-Sim-Cos@k=50 $88.06$ $82.09$ $94.46$
-Sim-Cos@k=100 $86.56$ $80.10$ $93.28$
-Sim-Cos@k=200 $75.12$ $78.23$ $86.76$
-Sim-Dot@k=1 $80.09$ $82.09$ $88.75$
-Sim-Dot@k=2 $81.34$ $83.45$ $90.05$
-Sim-Dot@k=5 $82.58$ $84.70$ $91.29$
-Sim-Dot@k=10 $84.57$ $85.32$ $92.60$
-Sim-Dot@k=20 $85.57$ $84.57$ $93.10$
-Sim-Dot@k=50 $86.07$ $81.96$ $93.59$
-Sim-Dot@k=100 $84.08$ $78.73$ $92.41$
-Sim-Dot@k=200 $75.12$ $77.73$ $86.76$
-Table 5: The effect of $k$ on detection performance with IMDB.
-Method 5% noise 10% noise 20% noise
-Sim-Cos@k=1 $28.50$ $50.85$ $66.15$
-Sim-Cos@k=2 $49.10$ $66.80$ $69.72$
-Sim-Cos@k=5 $57.80$ $72.05$ $75.65$
-Sim-Cos@k=10 $59.50$ $74.10$ $76.77$
-Sim-Cos@k=20 $59.50$ $74.90$ $77.42$
-Sim-Cos@k=50 $60.20$ $75.55$ $78.42$
-Sim-Cos@k=100 $59.90$ $75.50$ $78.52$
-Sim-Cos@k=200 $60.00$ $75.45$ $78.57$
-Sim-Dot@k=1 $57.50$ $67.25$ $75.17$
-Sim-Dot@k=2 $57.00$ $67.30$ $75.30$
-Sim-Dot@k=5 $57.20$ $67.95$ $75.80$
-Sim-Dot@k=10 $57.10$ $67.75$ $76.00$
-Sim-Dot@k=20 $57.10$ $68.15$ $76.27$
-Sim-Dot@k=50 $57.90$ $68.55$ $76.95$
-Sim-Dot@k=100 $59.90$ $68.60$ $77.55$
-Sim-Dot@k=200 $58.20$ $68.25$ $78.40$
-The effect of $\tau$ on error reduction rate.
-We analyze the effect of $\tau$ on the error reduction rate.
-We vary the number of $\tau=\{0.5,0.6,0.7,0.8,0.9,0.99\}$.
-From results in Tab.
-6 and Tab.
-7, we see that for both Snippets and IMDB, the change of $\tau$ has minimal impact on the error reduction rate.
-Generally, the reduction rate is higher when the noise level is higher.
-Table 6: Error Reduction Rate on Snippets.
-Method 5% noise 10% noise 20% noise
+3 shows the detection accuracy across different noise types and levels.
+Overall, Sim-Cos consistently outperforms Sim-Dot across all settings.
+However, both Sim-Cos and Sim-Dot achieve lower absolute accuracy with LLMs compared to the fine-tuned BERT model.
+This can be intuitively explained by the fact that, in a well-trained BERT, features from different classes tend to be pushed into different orthogonal subspaces (c.f.
+Section 3.2 and Csordás et al.
+).
+In contrast, LLMs are not task-specific trained, causing features of different classes to be more entangled, which reduces the effectiveness of similarity-based methods.
+6.  Conclusion
+We introduce similarity-based algorithms with theoretical justification for detecting and rectifying noise in datasets.
+Experiment results demonstrated the superior performance of our approach to improve the dataset quality and model generalization.
+Limitations
+We discuss the limitations of similarity-based methods: (1) The optimal detection accuracy of Sim-Cos and Sim-Dot, unfortunately based on empirical validation and depends on the choice of $k$ and $\mathcal{D}_{\textnormal{aux}}$, and also with different datasets and model architectures.
+(2) The generalization of models under the removal or rectification of noise remains uncertain, due to the limited exploration of datasets.
+Ethics Statement
+We consider only the public datasets and create artificial noises for evaluation.
+We do not pose any concern about the quality of the original datasets.
+AI tools were used for grammar checking, and technical content was written by the authors.
 […]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-Table 7: Error Reduction Rate on IMDB.
-Method 5% noise 10% noise 20% noise
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-B.3 The correlation between detection methods.
-We calculate the Spearman correlation between ranking scores assigned to samples by detection methods.
-Fig.
-9(a)a shows that gradient-based, confident-based, and similarity-based methods have low Spearman correlation with each other.
-Confident-based and Gradient-based methods have negative Spearman correlation, indicating they are very different in ranking.
-We observed the same phenomenon in datasets across levels and types of noise.
-(a) Visualization of (a) Spearman correlation, (b) the $L_{2}$-norm of noisy data points (presented by blue bars) and normal data points (presented by red bars).
-Using normalization (Cosine; (d)) enables the differentiation of the similarity distribution, therefore, enhancing detection accuracy compared to the unnormalized approach (Dot; (c)).
-Table 8: Test accuracy on Snippets with 10% noise.
-Method Random noise Ambiguity noise Concentrated noise
-Removed Rectified Removed Rectified Removed Rectified
-Acc. (under noise) 85.43 (+0.00) 88.24 (+0.00) 85.08 (+0.00)
-Confidence-W. Entropy 88.07 (+2.64) 85.04 (-3.20) 84.56 (-0.52)
-Normalize-Margin 81.75 (-3.68) 86.09 (-2.15) 79.73 (-5.35)
-Self-Confidence 84.47 (-0.96) 87.06 (-1.18) 82.06 (-3.02)
-Influence Function 90.04 (+4.61) 85.92 (-2.32) 81.36 (-3.72)
-Gradient-Cosine 82.94 (-2.49) 81.18 (-7.06) 80.48 (-4.60)
-Gradient-Dot 86.01 (+0.58) 85.13 (-3.11) 83.29 (-1.79)
-TracIn 85.74 (+0.31) 86.88 (-1.36) 77.98 (-7.10)
-Sim-Cos 79.73 (-5.70) 87.85 (+2.42) 85.48 (-2.76) 85.57 (-2.67) 81.58 (-3.23) 82.06 (-3.02)
-Sim-Dot 88.29 (+2.86) 87.32 (+1.89) 84.34 (-3.90) 86.58 (-1.66) 81.62 (-3.46) 80.83 (-4.25)
-Table 9: Test accuracy on Snippets with 5% noise.
-Method Random noise Ambiguity noise Concentrated noise
-Removed Rectified Removed Rectified Removed Rectified
-Acc. (under noise) 87.14 (+0.00) 87.36 (+0.00) 87.93 (+0.00)
-Confidence-W. Entropy 84.86 (-2.28) 86.57 (-0.79) 87.80 (-0.13)
-Normalize-Margin 84.65 (-2.49) 84.82 (-2.54) 83.77 (-4.16)
-Self-Confidence 85.35 (-1.79) 83.99 (-3.37) 86.09 (-1.84)
-Influence Function 85.30 (-1.84) 86.00 (-1.36) 85.30 (-2.63)
-Gradient-Cosine 85.43 (-1.71) 85.13 (-2.23) 79.51 (-8.42)
-Gradient-Dot 84.56 (-2.58) 89.21 (+1.85) 87.06 (-0.87)
-TracIn 84.65 (-2.49) 87.71 (+0.35) 81.88 (-6.05)
-Sim-Cos 86.45 (-0.69) 85.17 (-1.97) 87.10 (-0.26) 85.48 (-1.88) 85.39 (-2.54) 87.19 (-0.74)
-Sim-Dot 87.15 (+0.01) 87.06 (-0.08) 83.37 (-3.99) 84.34 (-3.02) 84.12 (-3.81) 79.82 (-8.11)
-Table 10: Test accuracy on IMDB with 20% noise.
-Method Random noise Ambiguity noise Concentrated noise
-Removed Rectified Removed Rectified Removed Rectified
-Acc. (under noise) 89.92 89.36 85.98
-Confidence-W. Entropy 91.00 (+1.08) 90.73 (+1.37) 86.66 (+0.68)
-Normalize-Margin 91.00 (+1.08) 90.73 (+1.37) 86.66 (+0.68)
-Self-Confidence 91.00 (+1.08) 90.73 (+1.37) 86.66 (+0.68)
-Influence Function 90.95 (+1.03) 90.51 (+1.15) 87.09 (+1.11)
-Gradient-Cosine 90.59 (+0.67) 90.93 (+1.57) 86.78 (+0.80)
-Gradient-Dot 90.43 (+0.51) 89.45 (+0.09) 82.87 (-3.20)
-TracIn 90.19 (+0.27) 91.09 (+1.73) 82.87 (-3.20)
-Sim-Cos 90.70 (+0.78) 88.73 (-1.19) 90.78 (+1.42) 90.83 (+1.47) 87.76 (+1.78) 84.73 (-1.25)
-Sim-Dot 91.49 (+1.57) 90.46 (+0.54) 91.58 (+2.22) 90.13 (+0.77) 87.25 (+1.27) 87.46 (+1.48)
-Table 11: Test accuracy on IMDB with 10% noise.
-Method Random noise Ambiguity noise Concentrated noise
-Removed Rectified Removed Rectified Removed Rectified
-Acc. (under noise) 91.76 (+0.00) 91.68 (+0.00) 89.41 (+0.00)
-Confidence-W. Entropy 91.79 (+0.03) 92.48 (+0.20) 90.95 (+1.54)
-Normalize-Margin 91.79 (+0.03) 92.48 (+0.20) 90.95 (+1.54)
-Self-Confidence 91.79 (+0.03) 92.48 (+0.20) 90.95 (+1.54)
-Influence Function 91.65 (-0.11) 90.97 (-0.71) 86.26 (-3.15)
-Gradient-Cosine 91.72 (-0.04) 91.29 (-0.39) 88.05 (-1.36)
-Gradient-Dot 90.81 (-0.95) 92.10 (+0.42) 88.15 (-1.26)
-TracIn 91.75 (-0.01) 91.66 (-0.02) 88.86 (-0.55)
-Sim-Cos 91.12 (-0.64) 91.98 (+0.22) 92.52 (+0.84) 92.17 (+0.49) 90.15 (+0.74) 89.20 (-0.21)
-Sim-Dot 91.38 (-0.38) 91.51 (-0.25) 91.52 (-0.16) 91.61 (-0.07) 89.78 (+0.37) 88.37 (-1.04)
-Table 12: Test accuracy on IMDB with 5% noise.
-Method Random noise Ambiguity noise Concentrated noise
-Removed Rectified Removed Rectified Removed Rectified
-Acc. (under noise) 91.67 (+0.00) 92.07 (+0.00) 91.52 (+0.00)
-Confidence-W. Entropy 92.35 (+0.68) 92.58 (+0.51) 91.94 (+0.42)
-Normalize-Margin 92.35 (+0.68) 92.58 (+0.51) 91.94 (+0.42)
-Self-Confidence 92.35 (+0.68) 92.58 (+0.51) 91.94 (+0.42)
-Influence Function 92.55 (+0.88) 91.76 (-0.31) 90.77 (-0.75)
-Gradient-Cosine 92.46 (+0.79) 90.88 (-1.19) 89.78 (-1.74)
-Gradient-Dot 91.79 (+0.12) 92.58 (+0.51) 90.32 (-1.20)
-TracIn 92.62 (+0.95) 91.97 (-0.10) 91.45 (-0.07)
-Sim-Cos 92.10 (+0.43) 92.10 (+0.43) 93.09 (+1.02) 92.50 (+0.43) 92.11 (+0.59) 91.74 (+0.22)
-Sim-Dot 92.69 (+1.02) 92.06 (+0.39) 91.63 (-0.44) 91.09 (-0.98) 91.59 (+0.07) 90.96 (-0.56)
-Figure 9: Detection accuracy of methods measured on Snippets with different levels of noise.
-Figure 10: Detection accuracy of methods measured on IMDB with different levels of noise.
+  • Agarwal et al.
+(2017) Naman Agarwal, Brian Bullins, and Elad Hazan.
+2017.
+Second-order stochastic optimization for machine learning in linear time.
+The Journal of Machine Learning Research, 18(1):4148–4187.
+  • AI@Meta (2024) AI@Meta. 2024. Llama 3 model card.
+  • Bahri et al. (2020) Dara Bahri, Heinrich Jiang, and Maya Gupta. 2020. Deep k-nn for noisy labels. In International Conference on Machine Learning, pages 540–550. PMLR.
+  • Basu et al. (2021) Samyadeep Basu, Phil Pope, and Soheil Feizi. 2021. Influence functions in deep learning are fragile. In International Conference on Learning Representations.
+  • Beyer et al. (2020) Lucas Beyer, Olivier J Hénaff, Alexander Kolesnikov, Xiaohua Zhai, and Aäron van den Oord. 2020. Are we done with imagenet? arXiv preprint arXiv:2006.07159.
+  • Brown et al.
+(2020a) Tom Brown, Benjamin Mann, Nick Ryder, Melanie Subbiah, Jared D Kaplan, Prafulla Dhariwal, Arvind Neelakantan, Pranav Shyam, Girish Sastry, Amanda Askell, Sandhini Agarwal, Ariel Herbert-Voss, Gretchen Krueger, Tom Henighan, Rewon Child, Aditya Ramesh, Daniel Ziegler, Jeffrey Wu, Clemens Winter, Chris Hesse, Mark Chen, Eric Sigler, Mateusz Litwin, Scott Gray, Benjamin Chess, Jack Clark, Christopher Berner, Sam McCandlish, Alec Radford, Ilya Sutskever, and Dario Amodei.
+2020a.
+Language models are few-shot learners.
+In Advances in Neural Information Processing Systems, volume 33, pages 1877–1901.
+Curran Associates, Inc.
+  • Brown et al.
+(2020b) Tom Brown, Benjamin Mann, Nick Ryder, Melanie Subbiah, Jared D Kaplan, Prafulla Dhariwal, Arvind Neelakantan, Pranav Shyam, Girish Sastry, Amanda Askell, et al.
+2020b.
+Language models are few-shot learners.
+Advances in neural information processing systems, 33:1877–1901.
+  • Charpiat et al.
+(2019a) Guillaume Charpiat, Nicolas Girard, Loris Felardos, and Yuliya Tarabalka.
+2019a.
+Input similarity from the neural network perspective.
+Advances in Neural Information Processing Systems, 32.
+  • Charpiat et al.
+(2019b) Guillaume Charpiat, Nicolas Girard, Loris Felardos, and Yuliya Tarabalka.
+2019b.
+Input similarity from the neural network perspective.
+Advances in Neural Information Processing Systems, 32.
+  • Chong et al.
+(2022) Derek Chong, Jenny Hong, and Christopher Manning.
+2022.
+Detecting label errors by using pre-trained language models.
+In Proceedings of the 2022 Conference on Empirical Methods in Natural Language Processing, pages 9074–9091, Abu Dhabi, United Arab Emirates.
+Association for Computational Linguistics.
+  • Chu et al.
+(2013) Xu Chu, Ihab F.
+Ilyas, and Paolo Papotti.
+2013.
+Holistic data cleaning: Putting violations into context.
+In 2013 IEEE 29th International Conference on Data Engineering (ICDE), pages 458–469.
+  • (12) Róbert Csordás, Sjoerd van Steenkiste, and Jürgen Schmidhuber.
+Are neural nets modular?
+inspecting functional modularity through differentiable weight masks.
+In International Conference on Learning Representations.
+  • Dau et al.
+(2022) Anh TV Dau, Nghi DQ Bui, Thang Nguyen-Duc, and Hoang Thanh-Tung.
+2022.
+Towards using data-influence methods to detect noisy samples in source code corpora.
+In 37th IEEE/ACM International Conference on Automated Software Engineering, pages 1–3.
+  • Devlin et al.
+(2019) Jacob Devlin, Ming-Wei Chang, Kenton Lee, and Kristina Toutanova.
+2019.
+BERT: Pre-training of deep bidirectional transformers for language understanding.
+In Proceedings of the 2019 Conference of the North American Chapter of the Association for Computational Linguistics: Human Language Technologies, Volume 1 (Long and Short Papers), pages 4171–4186, Minneapolis, Minnesota.
+Association for Computational Linguistics.
+  • Gao et al. (2016) Wei Gao, Bin-Bin Yang, and Zhi-Hua Zhou. 2016. On the resistance of nearest neighbor to random noisy labels. arXiv preprint arXiv:1607.07526.
+  • Goh and Mueller (2023) Hui Wen Goh and Jonas Mueller.
+2023.
+Activelab: Active learning with re-labeling by multiple annotators.
+In ICLR 2023 Workshop on Pitfalls of limited data and computation for Trustworthy ML.
+  • Goh et al.
+(2022) Hui Wen Goh, Ulyana Tkachenko, and Jonas Mueller.
+2022.
+Utilizing supervised models to infer consensus labels and their quality from data with multiple annotators.
+NeurIPS 2022 Human in the Loop Learning Workshop.
+  • Gunasekar et al.
+(2023) Suriya Gunasekar, Yi Zhang, Jyoti Aneja, Caio César Teodoro Mendes, Allie Del Giorno, Sivakanth Gopi, Mojan Javaheripi, Piero Kauffmann, Gustavo de Rosa, Olli Saarikivi, et al.
+2023.
+Textbooks are all you need.
+arXiv preprint arXiv:2306.11644.
+  • Hampel (1974) Frank R. Hampel. 1974. The influence curve and its role in robust estimation. Journal of the American Statistical Association, 69(346):383–393.
+  • Hanawa et al. (2021) Kazuaki Hanawa, Sho Yokoi, Satoshi Hara, and Kentaro Inui. 2021. Evaluation of similarity-based explanations. In International Conference on Learning Representations.
+  • Huang and He (2018) Zhipeng Huang and Yeye He. 2018. Auto-detect: Data-driven error detection in tables. In Proceedings of the 2018 International Conference on Management of Data, pages 1377–1392.
+  • Koh and Liang (2017) Pang Wei Koh and Percy Liang. 2017. Understanding black-box predictions via influence functions. In International conference on machine learning, pages 1885–1894. PMLR.
+  • Kuan and Mueller (2022) Johnson Kuan and Jonas Mueller. 2022. Model-agnostic label quality scoring to detect real-world label errors. In ICML DataPerf Workshop.
+  • Loshchilov and Hutter (2019) Ilya Loshchilov and Frank Hutter. 2019. Decoupled weight decay regularization. In International Conference on Learning Representations.
+  • Maas et al.
+(2011) Andrew L.
+Maas, Raymond E.
+Daly, Peter T.
+Pham, Dan Huang, Andrew Y.
+Ng, and Christopher Potts.
+2011.
+Learning word vectors for sentiment analysis.
+In Proceedings of the 49th Annual Meeting of the Association for Computational Linguistics: Human Language Technologies, pages 142–150, Portland, Oregon, USA.
+Association for Computational Linguistics.
+  • Nguyen-Duc et al.
+(2023) Thang Nguyen-Duc, Hoang Thanh-Tung, Quan Hung Tran, Dang Huu-Tien, Hieu Nguyen, Anh T.
+V.
+Dau, and Nghi Bui.
+2023.
+Class based influence functions for error detection.
+In Proceedings of the 61st Annual Meeting of the Association for Computational Linguistics (Volume 2: Short Papers), pages 1204–1218, Toronto, Canada.
+Association for Computational Linguistics.
+  • Northcutt et al.
+(2021a) Curtis Northcutt, Lu Jiang, and Isaac Chuang.
+2021a.
+Confident learning: Estimating uncertainty in dataset labels.
+Journal of Artificial Intelligence Research, 70:1373–1411.
+  • Northcutt et al.
+(2021b) Curtis G Northcutt, Anish Athalye, and Jonas Mueller.
+2021b.
+Pervasive label errors in test sets destabilize machine learning benchmarks.
+In Thirty-fifth Conference on Neural Information Processing Systems Datasets and Benchmarks Track (Round 1).
+  • Pezeshkpour et al.
+(2021) Pouya Pezeshkpour, Sarthak Jain, Byron Wallace, and Sameer Singh.
+2021.
+An empirical comparison of instance attribution methods for NLP.
+In Proceedings of the 2021 Conference of the North American Chapter of the Association for Computational Linguistics: Human Language Technologies, pages 967–975, Online.
+Association for Computational Linguistics.
+  • Phan et al.
+(2008) Xuan-Hieu Phan, Le-Minh Nguyen, and Susumu Horiguchi.
+2008.
+Learning to classify short and sparse text & web with hidden topics from large-scale data collections.
+In Proceedings of the 17th international conference on World Wide Web, pages 91–100.
+  • Pruthi et al.
+(2020) Garima Pruthi, Frederick Liu, Satyen Kale, and Mukund Sundararajan.
+2020.
+Estimating training data influence by tracing gradient descent.
+Advances in Neural Information Processing Systems, 33:19920–19930.
+  • Reeve and Kabán (2019) Henry Reeve and Ata Kabán.
+2019.
+Fast rates for a knn classifier robust to unknown asymmetric label noise.
+In International Conference on Machine Learning, pages 5401–5409.
+PMLR.
+  • Rolnick et al. (2017) David Rolnick, Andreas Veit, Serge Belongie, and Nir Shavit. 2017. Deep learning is robust to massive label noise. arXiv preprint arXiv:1705.10694.
+  • Srivastava et al.
+(2014) Nitish Srivastava, Geoffrey Hinton, Alex Krizhevsky, Ilya Sutskever, and Ruslan Salakhutdinov.
+2014.
+Dropout: A simple way to prevent neural networks from overfitting.
+Journal of Machine Learning Research, 15(56):1929–1958.
+  • Team et al.
+(2023) Gemini Team, Rohan Anil, Sebastian Borgeaud, Yonghui Wu, Jean-Baptiste Alayrac, Jiahui Yu, Radu Soricut, Johan Schalkwyk, Andrew M Dai, Anja Hauth, et al.
+2023.
+Gemini: a family of highly capable multimodal models.
+arXiv preprint arXiv:2312.11805.
+  • Thyagarajan et al.
+(2022) Aditya Thyagarajan, Elías Snorrason, Curtis Northcutt, and Jonas Mueller.
+2022.
+Identifying incorrect annotations in multi-label classification data.
+arXiv preprint arXiv:2211.13895.
+  • Touvron et al.
+(2023a) Hugo Touvron, Thibaut Lavril, Gautier Izacard, Xavier Martinet, Marie-Anne Lachaux, Timothée Lacroix, Baptiste Rozière, Naman Goyal, Eric Hambro, Faisal Azhar, et al.
+2023a.
+Llama: Open and efficient foundation language models.
+arXiv preprint arXiv:2302.13971.
+  • Touvron et al.
+(2023b) Hugo Touvron, Louis Martin, Kevin Stone, Peter Albert, Amjad Almahairi, Yasmine Babaei, Nikolay Bashlykov, Soumya Batra, Prajjwal Bhargava, Shruti Bhosale, et al.
+2023b.
+Llama 2: Open foundation and fine-tuned chat models.
+arXiv preprint arXiv:2307.09288.
+  • Wang et al.
+(2024) Haoyu Wang, Zhuo Huang, Zhiwei Lin, and Tongliang Liu.
+2024.
+NoiseGPT: Label noise detection and rectification through probability curvature.
+In The Thirty-eighth Annual Conference on Neural Information Processing Systems.
+  • Wang and Mueller (2022) Wei-Chen Wang and Jonas Mueller. 2022. Detecting label errors in token classification data. arXiv preprint arXiv:2210.03920.
+  • Wu et al.
+(2020) Pengxiang Wu, Songzhu Zheng, Mayank Goswami, Dimitris Metaxas, and Chao Chen.
+2020.
+A topological filter for learning with label noise.
+Advances in neural information processing systems, 33:21382–21393.
+  • Zhou et al.
+(2023) Chunting Zhou, Pengfei Liu, Puxin Xu, Srini Iyer, Jiao Sun, Yuning Mao, Xuezhe Ma, Avia Efrat, Ping Yu, Lili Yu, et al.
+2023.
+Lima: Less is more for alignment.
+arXiv preprint arXiv:2305.11206.